A Supplemental Bibliography of Publications about the Reduce Symbolic Algebra Language

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

27 April 2024 Version 2.32

Title word cross-reference

#11 [WR79]. #2 [Hea72b]. #3 [Fit73]. #4 [Fit73]. #8 [Har77a].

1 → 2 [Gro83]. 23 ≤ n ≤ 26 [Ng89c, Ng89b, Ng89a]. 3 [Eas91]. _{n} [Ng89c]. BŜO_{n} < 16 > [Ng89b]. BŜO_{n}⟨16⟩ [Ng89a]. ℓ^{∗} [KKV04a]. f(x, y) [CC88, CC89]. I [SMH98]. ↔ [Alv00]. N = 1 [KKV02, KKV03a, KKV04b, KVK03]. T · [Ano90]. X^{T} R [DÜ92b, DÜ92a].

-coverings [KKV04a]. -D [Eas91].

11th [IEE78]. 13th [HR92]. 1800 [GK80b]. 1st [Ano90].

2 [BOH78, BO80, CNxx, Dul87, Dul89, Ede81, Hea71, Hus81, Loo72, Ozixx, Rod84].

2. [Gro88a]. 2008 [KLS12]. 2009 [Hea09]. 286-based [YA89, YA87]. 2n [Hea72b].

3 [DMR88, DMR90, Eas87, Kaz87]. 3-D [Eas87]. 360 [Kal82].

4. [Gro88b]. 440 [Mel74]. 4th [DN93].

5 [Sag89]. 5. [Gro88b]. 5th [Ano88, NBC92].

6th [GT94, Mor89].

7. [Gro88c]. 74 [Hus81]. ’76 [Jen76]. ’79 [Ng79].

8. [Gro88c]. ’85 [BC85]. 86 [YA87]. 86/286 [YA87]. 86/286-based [YA87]. ’87 [Dav89]. ’88 [Ano88, Gia89, Mar93]. ’89 [Gon89].

’90 [WN90]. ’91 [Wat91]. ’92 [DN93, Fit93b]. ’94 [GT94]. ’96 [CL96].

AAECC [Mor89]. AAECC-6 [Mor89]. Abel [KLS12]. Abelian [KR86, KRxx]. Aberration [NSW85]. Absolute [Wan85, Cap86a]. Abstracts [Loo77]. AC [NR97]. AC-Theories [NR97]. Accelerator [AB89, AZZ91]. ACM [Gon89, Jen76]. ACM-SIGSAM [Gon89]. Adaptive [SAK^{+}88]. Adding [van86a]. Addition [Gro88b]. Adelaide [NBC92]. Adjoint [SV92]. Advanced [Vis73]. Affinity [BGH93]. aid [GV92]. aided [Ano90, BC91, DJ89, SMU^{+}89]. Algebra [Asl96, BGH93, BC85, Cap90, Dav88, Dav89, Div91, Eis90, EC87, FBC86, GL88, GL89, Grä95, Gra94b, Gri75, GK78b, Har89b, Hea82, HS95, MW91a, M^{+}94, Mor89, ON90, OT87, Rd91, WH89, YP91, BCDS87, BC91, Cal94, DZ00, DS97, EP91, FK89, GSZ85, Gra94a, GKW03, GK78a, HH06, Har89a, Hea73a, HPR96, Her12, IKRT89b, Kea92, Ken80, Oll88, PR02, PBG14, RLBS91, SMH98, SHWZ88, SHWZ89, SHWZ93, Ste94, Ueb92, Vul02]. Algebraic [ACM94, Ack85, BGK86, Boy93, Bro93, Dau79, DJ80, DJRR81, Gon89, Gri74, Hea68, Hea76, Jen76, Kan75, KR87, MW91a, M^{+}94, MW91b, Mel93b, MS79, Mor89, Nun95, PH81, Ray87, Ray90, Ric92, Sto77, SS91, SI87, UAYS87, Veg91, WN90, Wat91, BO80, CH85, EP90, Fit85, Gia89, Gon91, Hea71, Ng79, Sag88, Sag89, SMU^{+}89, Tao90, PH83]. Algebras [HT95, Sch91]. Algébrico [dos88]. Algorithm [BB93, GM88, Sch85b, Sny93, AM90, Bra90, FGPF89, IKRT89a, Kem81, Nor90]. Algorithms [Grä95, LS91, Mor89, SL92, SS91, WD85]. Ampère [KKV04d, VVKK12]. amplitudes [BL91]. Analitik [Hus81]. Analitik-74 [Hus81]. Analysis [Ack85, BCBBxx, Eas87, Eis90, EP91, GL88, GL89, Gaw89, Gre87, RT85, Sch83, Sto77, vH82, Aok89, BCBB90, Bru93, Eas91, ET89, GV92, GMS93, Oll88, Ren91, Ren92, SMU^{+}89]. Analytic [CNxx, Loo72, OT87, EKR89, US91]. analytical [Hie91, Rod84]. analyzing [GMS90]. and/or [Ren92]. Angular [Gro88b]. Annual [Ano88, IEE78]. Antitranslator [Kry84]. any [Wan85]. Anymore [Sim92]. ao [dos88]. Application [BO80, BC81, CNxx, CC88, Dau79, DZ00, Eis90, EC87, GL88, GL89, GMS90, GMS93, Ozixx, PH81, PR84, RLBS91, UYSA89, WH89, YP91, CC89, Dye94, FK89, GSZ85, KR93, TSD91]. Applications [DR86, ET89, NSW85, Som85, Spi87, GKW03, KKV04b, Kra03, NBC92, RG91, Spi89, SS91]. Applied [Dav88, FBC86, Mor89, SHM87, KR68]. Applying [MS79]. Approach [Cal72, FKM95, GKKV08, KKV03b, KKV04b, Mar93, Ren91]. Approximate [Nuc92, YK90]. Approximation [CC88, CC89, GS90]. Approximations [KRL86, KLR93]. April [BC85, Fit93b]. APS [GT94]. Arbitrary [Sas79, IKRT89b]. arising [KR93]. Arithmetic [Sas79]. Artificial [HR92]. Asilomar [IEE78]. Assisted [Gre87]. astro [DJ89]. astro-geophysics [DJ89]. Astrophysics [Dau79]. ATENSOR [IK96]. Atom [VR87]. August [DN93, GT94, Jen76, WN90]. Austria [BC85]. Automatic [BCBB90, Bar92, BBB92, BCBBxx, JL94, Mel93a, Sto77, Bat85, Gat85a]. Automatically [Sch85a]. automation [LD90]. Autonomous [Sch85b, Sch86, Sch88b]. Avalanches [Sav90, Sav88]. Axially [DÜ92a, DÜ92b].

B1700 [GK79, GK80b]. B1700/B1800 [GK79, GK80b]. B1800 [GK79, GK80b].

Backlund [GKZ91, FK89, Nuc90]. Baecklund [KF87]. Based [HS95, Ina80, UAYS87, Alv00, Boc92, GV92, Hv85, Nor90, Ren92, YA87, YA89]. Bases [BGK86, HT95, MMN88, van86b]. Bath [Fit93b]. be [Fit89]. beam [US91]. Beginners [WH89, SHWZ88, SHWZ89, SHWZ93]. benchmark [MH85]. bending [Bar92]. betatron [AZZ91]. between [Dav82]. Bibliography [ZZZxx]. Bifurcation [LP90]. BIGNUM [Gri81]. Bilinear [Ito88]. Binding [BGH93]. body [Oll88]. Bond [Gaw89]. Bonn [Wat91]. Boolean [Kon95]. Bounded [Ric92]. Boussinesq [KKV06]. brackets [KK93, KK95a]. Brazil [M^{+}94]. Brazilian [MW91a, M^{+}94]. Brussels [Ano88]. BSO [Ng89c]. Buchberger [Bra90, GM88]. Buckling [RT85, EP90, EP91]. Buffering [KC94]. Built [NR97]. Built-in [NR97]. Burroughs [GK79, GK80b].

C [Hie91, Ste94]. CAD [SAK^{+}88, SKA88]. Calculate [KF87]. Calculating [BGK86, Nuc90, Nuc92, AM90]. Calculation [Cal72, DR85, Dul87, Ede81, GKI02a, Gro83, MS79, PBG14, Vul03, van86b, BL91, Dul89, IKRT89a, JL94, Kem81, Gro88b]. Calculations [CB76, DMR88, ST89b, DMR90, IKRT89b, IKRT91, Lan94, Rod84, Sei91, SS91]. Cálculo [dos88]. Calculus [DÜ92a, Har89b, Sch82a, SHM87, ST89a, DÜ92b, Har89a, Kad96]. CALI [Gra94a, Gra94b]. calibration [BO80]. California [IEE78]. CAMAL [Fit93a]. Camassa [GKKV08]. Can [Fit89]. Cartan [HT91, KK06]. CAS [BGS93]. Case [BGH93, Hea82, WR79]. catastrophe [WD85]. Causal [Har96]. caustic [Ano90]. Caustics [Ioa90b, Ioa91b]. center [FGPF89]. centroids [Mat89]. Certain [Köl85, Mal82]. chains [PBG14]. chamber [BO80]. Champs [Cap89]. Characteristics [DR85]. Charmonium [Gro88c]. Chebyshev [Boy93]. Checking [MN92]. chiral [BL91, Lan94]. chromodynamics [Gro90b]. Clarifications [AE83]. Classical [Nuc90, Nuc92, KK00, Gro88a]. Classification [GZ90a, Zha93, GSZ85]. Classroom [Par14]. Clifford [BCDS87, HT95]. Closed [Man93]. Code [Gat76, Gat86, KK89, vanxx, Bat85, BGV94, Bru93, GV92, Gat85a, vHGv89]. coded [GK78b, GK78a]. Codes [Mor89]. cohomological [IKK03]. Cohomology [Ng89b, Ng89c, Ng89a]. collection [Tha89a, Tha89b]. college [Tha89a, Tha89b]. Collineation [MN92]. Collision [LD87, LD89]. Colloquium [Vis73]. column [YK90]. comment [Sch88a]. Comments [Ken80]. Commutative [Gra94b, Gra94a, Rod84]. commutator [Sei91]. Compact [Nor93, Nor95]. Compactified [DR85]. Comparison [Hus81, NW83]. Compilation [KRR87]. Complete [KK02b, Cap86a, KK06]. completion [Nor90]. Complex [Asl96, US91]. compression [Hv85]. Comput [Sag89]. Computation [ACM94, BHPS86, Bro93, DR86, GT91, Gon89, HI89, Hug90, Jen76, MW91b, MMN88, NM89, Ng89c, Ng89b, Ng89a, PSZ91, Ray87, Ray90, RT85, Sav90, UYSA89, WN90, Wat91, CL96, Dew89, Dye94, Fit93b, FGPF89, Gia89, LS96, Ng79, Sav88, SL92]. Computational [NBC92, HM24, KK95b]. Computations [Dau79, DJ80, DJRR81, KR86, KRxx]. Compute [Par14, Mat89]. Computer [Asl96, BGH93, BC91, BC85, Cap90, DMR88, Dav88, Dav89, Div91, DS97, DJ89, Edn93, FK89, GL88, GL89, GSZ85, GZ90a, GZ90b, GKW03, Gre87, Gri75, GK79, GK80b, HS95, HPR96, MW91a, M^{+}94, Nor80, OT87, Rd91, SHM87, SMH98, SHWZ88, SHWZ89, SHWZ93, Sto77, WH89, YP91, Zha93, Cal94, DMR90, DZ00, EKR89, EP91, Hea73a, Her12, IKRT89b, Kea92, Ken80, Oll88, PBG14, RLBS91, SMU^{+}89, Ste94, SS91, Ueb92, Vul02]. Computer-aided [DJ89, BC91]. Computer-Assisted [Gre87]. Computeralgebra [Ueb92]. Computerized [EP90, Ano90]. Computers [YA87, YA89]. Computing [BKR^{+}91, GT94, KRR88, Man93, Nun95, Ric92, Vis73, DN93, HPR96, HR92, MW91a, M^{+}94]. conditions [GZ90b, KKV04c]. Conference [Ano88, BC85, Dav89, GT94, IEE78, Mor89, NBC92, DN93]. Congress [HR92]. Congressi [GT94]. Conserved [IK85, GKI02a, GKI02b, Ito94]. consistency [GMS90, GMS93]. constants [YK90]. Constraints [NR97]. Construction [Ioa90b, Ioa90a, Ioa91b, Ioa91a, KRL86, PR84, LS91]. constructions [KKV04e]. Constructive [HH06, HT91, Ren91, Ren92]. Contact [EFK85]. Continuity [ST89a]. Control [SAK^{+}88, SKA88, UAYS87, UYSA89, ET89]. controlled [Hv85]. controller [LS96]. convection [PR02]. conversion [De 89]. CONVODE [MM93, Mou93]. Coordinates [Eas87, Har89b, Eas91, Har89a]. Correcting [Mor89]. corrections [Gro91a]. Correlation [PSZ91]. Coulomb [BH88]. Coupled [GZ90a, Zha93, GKZ91, KK02b]. covering [Kra03, KK06]. coverings [KKV04a]. CP [MF83]. CP/M [MF83]. Cracks [Ioa90a, Ioa91a]. Cray [AGK^{+}87]. critical [Nor90]. critical-pair [Nor90]. critical-pair/completion [Nor90]. Cross [Kot86]. Cross-Section [Kot86]. CSL [Tri00]. CSL-hosted [Tri00]. CTS [KR87]. Curvature [Gro88b]. Curve [Ric92]. curved [Vul03]. Curves [Veg91, Gon91]. Curvilinear [Eas87, Har89b, Eas91, Har89a]. Cycles [LP90]. cyclic [AZZ91]. Czechoslovakia [DN93].

D [Eas87, Eas91]. dans [Cap89]. Data [Gri76, PH83]. Data-Structures [Gri76]. Debugging [KR85a, KR87]. DEC [HS83]. decays [Gro83]. Decomposition [BCRS89]. Definite [Köl85]. Definition [Gri76]. deflection [Aok89]. Deformations [KKVV10, KK92, KK94, KK95b]. Delivery [Nor93, Nor95, Fit90]. Densities [IK85, GKI02b, Ito94]. derivation [DJ89]. Derivatives [Hv83, Kem81]. Description [Cap84, Cap86b, Gra94b, Gra94a]. Design [CL96, Fit93b, Gat85b, Hol92, UAYS87, UYSA89, ET89, LS96, Mar93, SMU^{+}89, YK90]. Determination [Sch91]. Determining [EFK85, KF87, Sch82b, Sch85b, Sch85a, Sch86, FK89, Sch88b]. Development [Hea82, Kry86a, Kry86b, SAK^{+}88, SMU^{+}89, HM24]. devices [BO80]. Diagnosing [Gre87]. diagram [Sch91]. Diagrams [Cal72, KR86, KRxx, Gro83, HM24]. Dialogue [Kry86a, Kry86b]. Diatomic [OT87]. Difference [GL88, GL89, GS90, GV92, GKI02a, GKI02b, LD90, LS91]. difference-difference [GKI02b]. Differential [EFK85, GK91, IK85, KF87, Mac88, Mal82, MM93, Nuc90, Nuc92, RT89, SV92, Sch82b, Sch85b, Sch85a, Sch86, BB93, Bil92, FK89, GMS93, GZ83, HT91, KK93, KK95a, KKV04c, KK95b, KK00, Mac89, Mou93, RT88, Sch88b, KLS12]. differentiation [JL94]. Differenzenoperationen [GF88]. Differenzenverfahren [GVZ91]. diffusion [YK90]. Digital [Hol92]. dimension [IKRT89b]. dimensional [PBG14]. Dimer [BGH93]. Dirac [IKRT89a, IKRT89b, IKRT91, Vul02, Vul03]. direct [Ano90]. DISCO [CL96, Fit93b]. dispersionless [KKV06]. Doing [HI89]. Domain [BHPS86]. Double [DÜ92a, DÜ92b]. drift [KLR93]. Duality [DÜ92a, DÜ92b]. Dublin [HR92]. dummy [IK91]. d’Une [Cap84]. Dutch [van88]. Dynamic [Ioa90b, KR85a, Ioa91b, LS96]. Dynamic-Debugging [KR85a]. Dynamics [FBC86]. Dynkin [Sch91].

Editing [KO83]. efficiency [Dav82]. Eigenparameter [Sha87]. Einführung [Ueb92]. Einstein [Sch88a]. Elastic [Eis90]. Elasticity [Ioa90b, Ioa91b]. Electric [VR87]. electrodynamics [Gro90c]. Element [BCBBxx, BCBB90, Bar92]. elementary [Gro90a, Gro90c, Gro90b, Gro91a, Gro91b]. Elements [BBB92, BH88]. Elliptical [Kot86]. Embedding [Rd91]. energy [BL91, De 89, Gro97, Gro05, SS91]. Engineering [SAK^{+}88, SKA88]. Enlarging [BHPS86]. Environment [Ede81, Dew89]. EPS [GT94]. EPS-APS [GT94]. EQSHELL [Boc92]. Equation [GK91, Ioa90b, Mac88, Mel93b, Mel93a, Sha87, GKKV08, Ioa91b, KKV02, KKV03a, KKV04d, KKV04b, KKV06, KVK03, Mac89, RT88, VVKK12, Vul03]. Equations [BGH93, BGK86, DÜ92a, EFK85, Gri74, Ioa90a, IK85, Kam05, KF87, KLS12, LD87, Mal82, McC84, MM93, Nuc90, Nuc92, RT89, SV92, Sch82b, Sch85b, Sch85a, Sch86, Sny93, BB93, Boc92, DZ00, DÜ92b, DJ89, FK89, GMS93, GKI02a, GKI02b, GZ83, GSZ85, GKZ91, IKK03, Ioa91a, Ito94, KR93, KK93, KK95a, KKV03b, KKV04c, KK92, KK94, KK95b, KK00, LD89, Loo72, Mou93, Ren91, Ren92, Sch88b]. Equilibria [Par14]. equilibrium [Boc92, KR93]. Erratum [Sag89]. Error [Ack85, Mor89, Sto77, Bam88]. Error-Correcting [Mor89]. ESPRIT [Ano88]. EUROCAL [BC85, Dav89]. European [BC85, Dav89]. EUROSAM [Ng79]. Evaluating [Ito85]. Evaluation [AB89, Har77b, Har79b, KRL93, Köl85, LD87, LD89, EP90]. Even [Maz85]. evolution [GSZ85, GZ90b, Ito94, KKV03b, KKV04c, KK92, KK94, Ren91]. evolutionary [Mar93]. evolutions [Ren92]. Exact [BH88, McC84, Kra03]. Examples [BGK86, FBC86]. EXCALC [SHM87, SMH98]. Exclusive [DMR88, DMR90]. Executing [KO83]. execution [PH83]. expansion [Hv85]. Expansions [GT91, PB90]. Experience [Kry86a, Kry86b, MS79]. experimental [KR68]. Experiments [AW89, van86b]. expert [HR92]. explanation [Her12]. Explicit [Köl85, Mal82, WD85]. expressed [YK90]. Expression [vH82, Hv85]. expressions [IK91]. Extended [KO83]. Extension [Cap84, Cap86b]. Extensions [ÜHK82, Spi87, Spi89]. extensive [KKV02]. Exterior [DÜ92a, Sch82a, SHM87, DÜ92b, HT91]. External [Dul87, Dul89].

Facilities [Hv83, MG88, vv88, vHv87]. Facility [Gat76, Gat86, Bat85, Gat85a]. Factorization [NSW85, Hv85, KK02a, NW83]. Families [RT89]. Fast [Dav82, IKRT89a, Sny93, SI87, IKRT91, JL94, Kem81]. Fermions [Hug90]. Feynman [Cal72, Dul87, Dul89, HM24, KR86, KRxx]. FIDE [LD90]. Field [Ack85, Dul87, Gro88c, McC84, ST89b, VR87, Dul89, Vul02]. Fields [Ede81]. Filters [Hol92]. Finding [IK85, Mal82, McC84, GKI02b, Ito94]. Finite [BCBBxx, BCBB90, Bar92, LD90]. First [Hea05, MW91a, M^{+}94, Man93, Sch85b, Sch86, Sch88b]. five [Oll88]. five-body [Oll88]. flat [IKK03]. Flow [Kot86, Sav90, YP91, Sav88]. flows [KR93]. Fluid [FBC86, LD87, YP91, LD89, PR02]. foer [Har80]. Form [Ito85, Man93, Ren92, WD85, BS98]. Formal [Hv83]. formation [Ano90]. Formelbehandling [Har80].

Formelmanipulation [Mel74]. formula [Hie91]. formulae [Tao90]. Formule [van88]. Fortran [Ina80, Sha87, BGV94, KK89, Kea92]. Fortran-Based [Ina80]. Forty [Hea05]. Foundation [Eis90]. foundations [GKW03]. fracture [Ano90]. frames [Har96]. France [Ng79]. Fredholm [Loo72]. free [Hea09]. French [Cap89]. Frölicher [KK93, KK95a]. Function [Cap90, PSZ91, Cap86a]. Functions [Asl96, BBB92, BCBBxx, Hea72b, Maz85, AM90, BCBB90].

Game [Par14]. gamma [IKRT89a]. gamma-matrices [IKRT89a]. Gauge [KR86, KRxx, McC84, SHM87, McC94]. GCUR [Gon91, Veg91]. GDR [Dav89]. General [Dau79, DJ80, DJRR81, Ozixx, SHM87, KKVV13, M^{+}94, McC94]. generality [Dav82]. Generalized [FK86, VVKK12]. Generating [Sny93]. Generation [BBB92, BCBBxx, Edn93, Gat76, Gat86, BCBB90, Bar92, Bat85, Boc92, BGV94, GV92, Gat85a, GZ90b]. Gentran [Gat85b, Gat85c, Bat85, BGV94, Gat85a, Gat87]. GENTRANS [Kea92, Kea91]. geodesy [TSD91]. geometric [KKV03b, KKV06]. Geometrical [DR85, GKKV08]. Geometries [DÜ92a, DÜ92b]. Geometry [GK91, KLS12, KKV04e, Tha89a, Tha89b]. Geophysical [FBC86]. geophysics [DJ89]. German [Ueb92]. Germany [Cal94, CL96, Wat91]. gimbal [KLR93]. GITA [BCM^{+}95]. Gluonic [Dul87, Dul89]. Graded [GK91, KK93, KK95a, KK95b]. Granular [Sav90, Sav88]. Graph [Gaw89]. Graphical [LS85]. Graphics [Mar88, Mar89]. Graphs [Dul87, Dul89]. Grassmann [HT95]. gravitation [DZ00]. Gravity [Vul02, SMH98]. Gröbner [MMN88, BGK86, HT95, van86b]. Ground [IEE78, Sny93]. Group [Cap90, KR86, KRxx]. Group-Theoretical [KR86, KRxx]. Grove [IEE78]. Guide [DJRR81, HS83]. gyroscope [KLR93].

Hadamard [Sch83]. Hamiltonian [KK02a, KKV03b, KKV03a, KKV04a, KKV04d, KKV04b, KKVV13, KVK03, Ren92]. Hamiltonians [BCM^{+}95]. handbook [GKW03]. Harmonic [BCRS89, DJ89]. Heights [Jen76]. held [NBC92]. Help [Ioa90b, Ioa91b]. Heuristic [Mon92]. Hidden [Sha87]. hierarchies [KKV03a, KKV04d, KKV04b]. High [IKRT89b, Gro97, Gro05, SS91]. highest [KLR93]. Hilfe [GVZ91]. Hirota’s [Ito88]. HLISP [Kan75]. Holm [GKKV08]. hosted [Tri00]. hot [YK90]. hot-wire [YK90]. Hydrogen [VR87]. hypergeometric [AM90].

iAPX [YA87]. iAPX86 [YA89].

iAPX86/286 [YA89]. iAPX86/286-based [YA89]. Idempotents [Kon95]. identities [IK91]. II [Hus81]. IMACS [HR92]. Implementation [FF82, Gat85b, GK79, Kan75, RG91, Sch82a, BB93, CL96, Fit93b, HT91, IKRT91]. Implementing [AGK^{+}87, Fit83]. Improved [Maz85]. Improvement [MG88]. includes [Ste94]. indeterminateness [Kem81]. Indicating [Wan85]. indices [IK91]. Industrial [BC81]. Infeasible [Gre87]. Infinite [BBB92]. Instability [EC87]. Installation [GM88, HS83]. Integrability [KKVV10, KLS12, GZ90b, GKKV08, Har96, KK02b, KKV04c, KK06]. Integrable [GZ90a, VVKK12, Zha93, GSZ85]. Integral [Ioa90a, Sha87, Ano90, Ioa91a, Loo72]. Integrals [AB89, Köl85, LD87, Sch85b, Sch86, AM90, BS98, LD89, Sch88b]. integrated [Dew89]. integrating [DZ00]. Integration [CNxx, Har79a, KO83, Bil92, Nor90, SS83]. intelligence [HR92]. Interacting [Ioa90a, Ioa91a]. Interaction [BH88]. interactions [Gro91b]. Interactive [Hea68, KR68, KO83, KR85b, KR85c, LS85, Nuc90, Nuc92, ST89a]. interactively [Kal82]. Interface [ASW89b, LS85, Mar88, ASW89a, Mar89]. Interfacing [NF96]. International [ACM94, Bro93, CL96, Fit93b, Gia89, Gon89, GT94, Mor89, Ng79, NBC92, Vis73, WN90, Wat91, DN93]. interpolation [Tao90]. Introducão [dos88]. Introduction [BGS93, Gro90a, HI89, BC91, Ueb92]. invariants [GKKV08, IKK03]. investigation [AZZ91, RLBS91]. Involving [Köl85]. Ireland [HR92]. IRENA [Dew89, KR93]. Irreducibility [Mon92]. Isovector [Ede81]. ISSAC [Gia89, Gon89, WN90, Wat91]. ISSAC’93 [Bro93]. ISSAC’94 [ACM94]. Italy [Gia89]. Iterations [Cap90]. iterative [WD85]. Itself [Dav88]. IZIC [FKM95].

J [Sag89]. J. [Ray90]. Jacobian [JL94]. Janeiro [M^{+}94]. January [Hea09]. Japan [WN90]. Java [NF96]. Joint [GT94]. July [Bro93, Gia89, Gon89, HR92, NBC92, Wat91]. June [Dav89, Ng79, Vis73]. Just [Sim92].

Kähler [HT91]. Karlsruhe [Cal94, CL96]. KdV [GZ90a, GK91, KK02b, KKV02, KKV03a, KKV04b, KVK03, KK06, Zha93]. KdV-like [GZ90a, Zha93]. Keener [Ray90]. Kiev [Bro93]. kind [Loo72]. Konstruktion [GF88]. Kuhn [Kam05]. Kupershmidt [KKVV10].

Laminar [Kot86]. Language [BOH78, Gro88a, Kan75, Kry84, SI87, UAYS87, EKR89, Hea71, SMU^{+}89]. Languages [Boy93]. Laplace [Kaz87, Kaz89, Nor80]. Large [Aok89, Gri74]. Lax [Ito85]. lecture [MW91a, M^{+}94]. Lectures [WH89, SHWZ88, SHWZ89, SHWZ93]. Legendre [Boy93]. Leipzig [Dav89]. Library [ZZZxx, BKR^{+}91]. Lie [EFK85, FK89, GKZ91, KF87, Nuc90, Sch91, Sch82b]. Lie-Backlund [GKZ91]. Lie-Baecklund [KF87]. like [GZ90a, Zha93]. LILK [Bru93]. Limit [Har77b, Har79b, LP90, Kem81]. Linear [Gre87, Gri74, ON90, UYSA89, BB93, GSZ85]. Links [BKR^{+}91]. Linz [BC85]. LISP [BC91, FF82, GK79, GK80b, HM^{+}83, Ina80, AGK^{+}87, Kal82, MH85]. Lisp/360 [Kal82]. loads [EP90]. Local [GS90, Grä95, KKV03a, KKV04b]. Logarithms [Köl85]. logic [DS97]. logically [LS91]. long [Kra03]. loop [BS98, Gro83]. LOOPS [ST89b]. Low [PR02, BL91]. low-energy [BL91]. Lucky [Grä93]. Lugano [GT94].

M [MF83]. m.b.v [van88]. machine [SS83]. MACSYMA [Fit73, Kea91, Kea92, Tha89a, Tha89b]. Magnets [Ack85]. Magnus [KLR93]. Main [Zha93]. Mainframes [Sim92]. man [SS83]. manifolds [FGPF89]. Manipulatie [van88]. Manipulation [Boy93, FK86, Kad93, Kan75, Ng79, SKA88, Sto77, UAYS87, Wan85, Ano90, Aok89, BO80, ET89, Hea71, Mat89, SMU^{+}89, Tao90]. manipulators [SMU^{+}89]. Manual [Gat85c, Gat87, Hea67, Hea73b, Hea87, Hea91, vanxx, Hea74, HM^{+}83]. MAPLE [Ste94, Vul02, Vul03]. Mapping [BBB92]. March [Cal94]. Marseille [Ng79, Vis73]. Matching [Kes79]. Math [Sim92]. MATHEMATICA [Ste94, KC94]. Mathematical [DR86, FKM95, TSD91]. Mathematics [FBC86, BC91, KR68, RG91]. MathML [Alv00]. Matlab [PR02]. Matrices [Kon95, AW89, Bar92, IKRT89a, JL94]. Matrix [BH88, Cap86b, Zha93, IKRT91]. Means [KRL86, GS90, IKRT89b]. measurement [Ano90]. Measurements [Ack85]. Mechanical [KRL93]. Mechanics [BC81, NM89, Ste94]. meets [DS97, Fit93a]. meshes [LS91]. meson [Lan94]. META [Kes79, Mar78]. META/REDUCE [Mar78, Kes79]. method [Ano90, Aok89, Bar92, DZ00, EP90, JL94, KK02a, LD90]. Methods [Boy93, Vis73, HM24, SS91]. metrics [Sch88a]. Micro [GK78b, GK79, IEE78, GK78a]. Micro-coded [GK78b, GK78a]. Micro-programmed [GK79]. Microprocessor [Fit83]. Microprogramming [IEE78]. mKdV [KK02b, KK06]. Mode [Mel14, PH81, AE83, CH85, LS96, PH83]. Model [Hea76, BL91]. modelled [PR02]. Modelling [Div91]. models [Lan94]. Modified [DÜ92a, GK91, DÜ92b]. Molecules [OT87]. Moments [PSZ91, Mat89]. Momentum [Gro88b]. monde [Cap89]. Monge [KKV04d, VVKK12]. Monogenic [BCRS89]. motion [KLR93, RLBS91]. movie [Hie91]. moving [Har96]. Multidimensional [DR85]. multilinear [RG91]. Multiloop [ST89b]. Multiple [Asl96]. Multiple-Valued [Asl96]. Multivariable [UYSA89].

NAG [BKR^{+}91]. Nash [Par14]. necessary [GZ90b]. Network [ZZZxx]. networks [KR93]. Newman [DZ00]. newsletter [Loo77]. Nijenhuis [KK93, KK95a]. Ninth [NSW85]. Ninth-order [NSW85]. no [Sag89]. Non [KKV04b, Nuc90, Nuc92, Ric92, GSZ85, KKV03a, LS91, Rod84, SMH98, KR86, KRxx]. non- [SMH98]. Non-Abelian [KR86, KRxx]. Non-Algebraic [Ric92]. Non-Classical [Nuc90, Nuc92]. non-commutative [Rod84]. non-linear [GSZ85]. non-orthogonal [LS91]. Nonconservative [EC87]. Nonlinear [Edn93, Gro88a, Gro88c, KRL93, Mel93b, Mel93a, Ren91, AZZ91, Ano90, GKI02b, GZ90b, GKZ91, Ito94, KR93, Ren92, Gro88b]. Nonlocal [KKV04e, KVK03]. Nonstationary [EH88]. normal [Ren92, WD85]. normalization [BCM^{+}95]. Normalizing [Edn93]. Note [Nor80, Sch88a]. notes [MW91a, M^{+}94]. November [Ano88, IEE78]. number [PR02]. Numeric [BKR^{+}91, Gri77]. Numerical [Gat76, Gat86, Ano90, Dew89]. numerics [HPR96].

obtaining [KLR93]. Occurring [AB89]. Odd [Maz85]. ODE [BGS93, Edn93]. ODEs [Man93]. off [Dav82]. one [Gro83, PBG14]. one-dimensional [PBG14]. one-loop [Gro83]. OpenMath [Alv00, Tri00]. Operations [Ito88, GS90]. Operator [Ito88]. operators [CH85, KK93, KK95a, KK02a, KKV04a, KK92, KK94, KK95b, KK00, RG91]. optical [Ano90]. Optics [NSW85]. Optimisation [PH81]. Optimization [vanxx, vHGv89]. Optimizing [AGK^{+}87]. Option [Kal82]. Orbit [AB89]. orbital [RLBS91]. Order [HS95, Man93, SV92, BB93, NSW85]. Order-Sorted [HS95]. Ordinary [Mac88, Mal82, Sch82b, Sch85b, Sch86, BB93, GZ83, Mac89, Sch88b]. Oregon [Gon89]. Orientational [PSZ91]. Oriented [Hea68]. Orthogonal [Eas87, Har89b, RT89, Eas91, Har89a, LS91]. Orthogonalization [BOH78]. ORTHOVEC [Eas87, Eas91]. oscillations [AZZ91]. Oscillator [Gro88a, Gro88c].

P [Ray90]. Pacific [IEE78]. Package

[Ack85, BH88, DJ80, DÜ92a, Gra94b, Gri81, Kad93, MM93, Sas79, SHM87, Sch82b, Sch83, Sch86, UAYS87, vH82, vanxx, Bam88, BGV94, Cap86a, Cap86b, DÜ92b, GKI02b, GZ83, Gon91, Gra94a, Hv85, Ito94, Kad96, LD90, Mou93, Ren92, Sch88b, Sei91, Veg91, vHGv89]. packages [NW83]. Painlevé [Ren91, Ren92]. Pair [Ito85]. pair/completion [Nor90]. Palazzo [GT94]. papers [HR92]. paradox [Her12]. parallel [Fit89]. Parallelism [Wan85]. Parallelization [Bra90]. parameter [Ano90]. Parameters [Wan85, YK90]. Paramodulation [NR97]. Part [Tha89a, Tha89b]. Partial [IK85, Sch82b, Sch85a]. particle [Gro90a, Gro90c, Gro90b, Gro91a, Gro91b]. particular [Oll88]. Pattern [Kes79, KRR87, Ano90]. PC [HI89, IKRT91]. PDE [KKV04e, LD90]. PDE’s [GMS90, KKVV13]. Penrose [DZ00]. Personal [YA87, YA89]. Perturbation [CB76, Aok89]. perturbative [SS91]. phonon [PBG14]. Physical [Gro88a, Gro88b, Gro88c, Wan85, Dye94, SS91]. Physics [CNxx, DN93, GT94, Vis73, Gro90a, Gro90c, Gro90b, Gro91a, Gro91b, Gro97, Gro05, Ken80, SS91]. picture [BO80]. Pipe [Kot86]. Plane [Ioa90b, Ric92, Veg91, Gon91, Ioa91b]. Plasma [LD87, LD89]. Plates [Eis90]. PMETA [Kes79]. Poincaré [Her12, McC84, McC94, SHM87]. Point [EFK85]. points [Kem81]. Polynomial [RT88, Sch85b, BCM^{+}95, GZ90b]. polynomial-nonlinear [GZ90b]. Polynomials [BOH78, BCRS89, Mon92, Bil92]. Portland [Gon89]. possessing [IKK03]. possible [Vul03]. Power [FK86, PB90, JL94]. Powers [Köl85]. Prague [DN93]. Prandtl [PR02]. Precision [Sas79]. Preliminary [Hus81]. Prelle [Man93]. Preprocessor [KK89]. Pretty [Hv83, vv88, vHv87]. Pretty-Print [vHv87]. Primer [Mel14]. Primes [Grä93]. Print [vv88, vHv87]. Printing [Hv83]. Problem [CB76, Har77a, Hea72b, Nor78, WR79, YP91, Oll88, TSD91]. Problems [Fit73, GK80a, Gro88a, Gro88b, Gro88c, Ioa90b, KC94, OT87, VR87, BO80, Fit85, Ioa91b, US91]. Procedure [Man93].

Procedures [FK86, SV92, ST89b].

Proceedings

[ACM94, BC85, Gon89, GT94, Mor89, Ano88, Bro93, CL96, Dav89, DN93, Fit93b, Gia89, IEE78, Jen76, NBC92, WN90, Wat91]. process [Ano90, BL91]. Processes [DMR88, MS79, DMR90]. Processing [SI87, BO80]. Produced [KK89, Kea92]. Program [BOH78, CNxx, DJ80, DR85, Eas87, EFK85, GT91, Har89b, Har77b, Har79b, Ito85, IK85, Ito88, KF87, Mag81, Mal82, WR79, BCM^{+}95, Bil92, Boc92, Eas91, EKR89, Har89a, Hea73a, IK96, Mar93, Mat89]. Programme [Cap84]. programmed [GK79]. Programming [KRR88]. Programs [Ede81, Gre87, KR85a, KR87, Nuc90, Nuc92, PH81, PH83, SMH98, Ste94, SS91, Tha89a, Tha89b, US91]. Prolog [Gaw89]. Properties [ST89a]. Proposal [MG88, ÜHK82]. Prover [ST89a]. Provisional [FF82]. Puiseux [GT91]. pure [BC91]. putting [Ano88].

QCD [BL91, Dul87, Dul89, MS79]. QD [Sch83]. QED [MS79]. Quadratic [McC84]. quadrature [Tao90]. Quadrupole [Ack85]. quadtree [AW89]. quantity [GKI02a]. Quantum [Gro90c, Gro90b, Ste94, ST89b, Sag88, Sag89, Gro88c].

Radiative [Gro91a, Gro88c]. random [EH88]. Rank [IK85]. Rational [KRL86, GKI02a]. Real [Sas79, Veg91, Gon91]. Realization [KO83, AM90]. Realizations [DMR88, DMR90]. Really [KO83]. rectangular [LS91]. Recurrence [PR84, RT88]. recursion [KK93, KK95a, KK02a, KK92, KK94, KK95b, KK00, KVK03]. recursions [KKV03a, KKV04d, KKV04b]. Recursive [Kem81]. REDLOG [DS97]. REDUCE [Loo77, Mel74, Sag89, US91, WD85, AE83, AM90, ZZZxx, ASW89b, Aok89, BS98, BOH78, BO80, Bam88, BCBB90, Bar92, BBB92, BCBBxx, BCM^{+}95, Bat85, BL91, BB93, Bil92, Boc92, BH88, BC81, BGV94, BCDS87, BC91, BHPS86, Bru93, Cal72, Cap84, CH85, Cap86a, Cap89, CNxx, CC88, CC89, CB76, DMR88, DMR90, Dau79, DJ80, DJRR81, Dav82, DR85, DZ00, DÜ92a, Dul87, Dul89, Eas87, Eas91, Ede81, EKR89, EFK85, EH88, Fit73, Fit83, Fit85, Fit89, Fit90, Fit93a, FBC86, GL89, GMS90, GS90, GV92, GMS93, GKI02a, GKI02b, Gat76, Gat85a, Gat85b, Gat85c, Gat86, Gat87, GZ83, Gon91, Gra94a, Gra94b, GK91, GK80a, Gri74, Gri75, Gri76, Gri77, GK78b, GK78a, GK79, GK80b, Gro83, Gro88a, Gro90a, Gro90c, Gro90b, Gro91a, Gro91b]. REDUCE [Gro97, Gro05, Har80, Har89b, Har89a, Har77a, Har77b, Har79a, Har79b, Hea67, Hea68, Hea69, Hea71, Hea72a, Hea72b, Hea73a, Hea73b, Hea74, Hea76, HS83, HM^{+}83, Hea87, Hea91, Hea05, Hea09, Her12, Hie91, HI89, Hol92, HvH^{+}83, Hv83, Hv85, Hus81, IK96, Ina80, Ioa90b, Ioa90a, Ito85, IK85, Ito88, Ito94, Kad93, Kad96, Kal82, Kan75, KK89, Kaz87, Kaz89, Kea91, Kea92, Kes79, KF87, KO83, Kry86a, KR85a, KR85b, KRL86, Kry86b, KR87, KRR87, KRR88, KRxx, KR85c, Lan94, LS85, LD87, LD89, LD90, LP90, Mac88, Mac89, MW91a, Mag81, Mar78, MF83, MH85, Mar88, Mar89, Maz85, MW91b, McC84, McC94, MMN88, Mel93b, Mel93a, Mel14, MN92, MG88, MM93, Mou93, NM89, Ng89c, Ng89b, Ng89a, Nor78, Nor93, Nor95, NF96, Nuc90]. REDUCE [Nuc92, Nun95, ON90, Ozixx, PB90, PR02, PH81, PH83, PR84, Ray87, Ray90, Ren91, Ren92, RG91, RT89, Sag88, SAK^{+}88, SV92, Sas79, Sch82a, SHM87, Sch82b, Sch83, Sch86, Sch88b, Sei91, Sha87, Som85, Spi87, Spi89, SL92, Ste94, ST89b, Tha89a, Tha89b, TSD91, ÜHK82, Ueb92, Veg91, VR87, Vul02, Vul03, Wan85, WR79, WH89, YA87, YA89, dos88, dos89, vv88, vH82, van88, vHGv89, vanxx, van86a, van86b, vHv87, Ack85, Alv00, Cap86b, De 89, DÜ92b, FK86, GL88, GF88, GVZ91, Gaw89, GT91, Har96, Hea82, Ioa91b, Ioa91a, Kam05, KR86, Loo72, Mal82, MS79, NW83, NV14, Par14, PBG14, Rod84, SKA88, SMU^{+}89, SMH98, SI87, UAYS87, ASW89a, Tri00]. REDUCE-2 [BOH78, BO80, CNxx, Dul87, Dul89, Ede81, Hus81, Rod84]. REDUCE-3 [DMR88, DMR90]. REDUCE-based [Boc92, Alv00]. REDUCE-Package [SHM87]. REDUCE-Procedures [SV92]. Reduce-programs [US91]. Reduce-system [GF88]. Reduce-systems [GVZ91]. REDUCE. [Gro88a, Gro88b, Gro88c]. REDUCE/ [GK78a]. REDUCE/1700 [GK78b]. REDUCE/1800 [GK80b]. Reduced [Kon95, Sny93]. Related [van86b, KKV04c]. relation [RT88]. Relations [PR84]. Relativity [Dau79, DJ80, DJRR81, HPR96, Ozixx, SHM87, M^{+}94, McC94]. Renormalization [Cap90]. Report [Sch82a, Ano90, HM^{+}83]. representation [AW89]. Representations [DR86, IKK03]. resonance [AZZ91]. restricted [Oll88]. Resultant [GV92]. Resultantenalgebra [GVZ91]. reuse [Mar93]. Review [Ber96]. revised [HR92]. Rewriting [Sny93]. Rhine [Cal94]. Ricci [Kad96]. Riemannian [DR85, SMH98]. RIKEN [Som85]. Rio [M^{+}94]. RLISP [Kry84, Mar93]. robot [SMU^{+}89]. Rome [Gia89]. root [Sch91]. rotational [OT87]. Rotator [Gro88c]. Rule [Mag81]. Rules [KRR88]. run [Fit89]. Running [HS83, Kal82].

sample [Ste94]. satellite [RLBS91]. Satisfying [DÜ92a, DÜ92b]. Scattering [MS79]. Schemes [GL88, GL89, Sch83, GV92, LS91]. School [MW91a, M^{+}94]. Schrodinger [GKZ91]. sciences [Dye94]. Scientific [DR86, HPR96]. SCOPE [vanxx]. Second [SV92, BB93, Loo72]. Second-Order [SV92]. Section [Kot86]. selected [HR92]. Semigroup [Kon95]. semisimple [Sch91]. September [CL96]. sequence [Kra03]. Sequencing [Mag81]. Series [FK86, HS83, PB90, Sch83]. Server [FKM95]. Set [Sny93]. Shape [BCBBxx, BCBB90]. Sharing [Kal82]. shells [Boc92]. Short [Gra94b, Gra94a]. Significance [Wan85]. SIGSAM [Gon89, Nor78]. similar [Loo72]. Simple [vHv87]. Simplification [Hea68, Hea76, Maz85, Cap86a, IK91, IK96]. simulation [Ano90, JL94, SHWZ88, SHWZ89, SHWZ93]. Singer [Man93]. Singular [Ioa90a, Ioa91a]. Sistema [dos88]. Six [WH89]. sliding [LS96]. Small [KO83]. Snitko [EP90]. Software [Ack85, KC94, Ray87, Ray90, Sim92, UAYS87, Hea09]. solar [De 89]. Solid [CNxx]. Solution [Gri74, Har77a, Hea72b, Mel93b, Mel93a, OT87, SI87, GZ83, Nor78, RT88, US91]. Solutions [Mal82, Man93, McC84]. Solve [Kam05]. Solved [WR79]. Solver [Mac88, Sha87, Mac89]. Solvers [Ber96]. Solving [BGK86, Fit85, Gro88a, Gro88b, Gro88c, KC94, BB93, LD90, Mou93, TSD91]. Some [BGK86, BKR^{+}91, DMR88, DMR90, Hv83, MS79, OT87, RT89, Spi87, Spi89, van86b, vHv87, KKV04c, US91]. Sorted [HS95]. Source [KK89, vanxx]. Source-Code [vanxx]. Space [DR85, DÜ92a, Wan85, DÜ92b, IKRT89b, SMH98]. Space-time [DÜ92a, DÜ92b]. space-times [SMH98]. Spacetime [Rd91]. spacetimes [Vul03]. Sparse [Gri74]. spectra [PBG14]. Spectral [Boy93, DJ89]. Spectroscopy [OT87]. speed [IKRT89b]. spherical [DJ89]. Spin [CB76]. Sprak [Har80]. Stabilitätsuntersuchung [GVZ91]. Stability [GL88, GL89, SI87, GV92]. Standard [FF82, GK80b, HM^{+}83, GK79].

State [CNxx]. statements [van86a]. static [Bru93]. statics [Mat89]. Stationary [DÜ92a, DÜ92b]. Status [Sch82a]. Steady [Kot86]. Straight [Ioa90a, Ioa91a]. Stratified [YP91]. Structural [NM89]. structure [Har96]. Structures [DR86, GK91, Gri76, KK02a, KKV03b, KKV03a, KKV04c, KKV04d, KKV04b, KKVV13, KVK03, PH83, VVKK12]. Study [Ano90, DÜ92a, Hea82, SV92, WR79, DÜ92b, GS90, KKV02, KKV06, KLR93]. Subroutines [DJRR81, Kea92]. Subscripts [Hv83]. Substitution [KRR88]. super [KK93, KK95a]. SUPERCALC [Sei91]. Supercomputer [MMN88]. Supergravity [GK80a]. supersymmetric [KKV02, KKV03a, KKV04b, KK00, KVK03]. Supersymmetry [GK91, GK80a, dos89]. Support [Nor93, Nor95, Tri00]. Surfaces [FKM95]. suspension [KLR93]. Switzerland [GT94]. symbol [Mat89]. Symbolic [ACM94, AB89, Ber96, Bro93, BKR^{+}91, DR86, Eis90, EC87, FBC86, GT91, Gia89, Gon89, Gri77, Har77b, Har79a, Har79b, HI89, Hug90, IK91, Jen76, KRL93, LS96, Mel93a, Mel14, NM89, Ng79, NR97, PSZ91, RT85, Sag89, SKA88, Sav88, Sav90, Sim92, Tao90, UAYS87, UYSA89, WN90, Wat91, AE83, Ano90, Aok89, CL96, Dew89, Dye94, ET89, EP90, Fit93b, FGPF89, HR92, Ken80, PR02]. Symbolic-Numeric [Gri77]. SYMCD [Ito94]. Symmetric [DÜ92a, DÜ92b]. Symmetries [EFK85, KF87, KK00, KLS12, Nuc90, Nuc92, SV92, Sch82b, Sch85a, FK89, GKZ91, IKK03, IK91, Ito94]. symplectic [KKV03a, KKV04d, KKV04b, KVK03]. Symposium [ACM94, Bro93, CL96, Fit93b, Gia89, Gon89, Jen76, KLS12, Ng79, WN90, Wat91]. Symsac [Jen76]. System [Dul87, EC87, GL88, GL89, Gri75, GK78b, Har79a, Hea68, Hea82, HS95, Ina80, KO83, KR85a, KR87, KRR88, Mar78, MS79, SAK^{+}88, Spi87, ST89b, VR87, AM90, Aok89, BO80, Dul89, ET89, Fit90, GMS90, GS90, GV92, GMS93, GK78a, Hea71, Her12, IKRT89b, JL94, KK02b, KK06, PBG14, Rod84, SMU^{+}89, Sch91, Spi89, TSD91, GF88]. Systematic [Ack85]. Systems [BGK86, Edn93, GZ90a, Gri74, KRL93, Mel93b, Mel93a, Sch85b, Sch86, Sny93, UYSA89, Zha93, CL96, Fit93b, GMS90, GMS93, GKI02b, GZ90b, GKW03, HH06, HT91, HR92, Ito94, Kea92, KR93, KR68, Sch88b, US91, GVZ91].

Tables [Nor80]. Tandem [Sha87]. Techniques [BGH93, PH81, NBC92]. technology [Ano88]. Tensor [Gro88b, Kad93, IK91, IK96]. terms [YK90]. Test [Mon92]. TEX-REDUCE [ASW89b, ASW89a]. their [KK95b]. Theorem [NSW85, Sch83, ST89a]. Theorem-Prover [ST89a]. Theoretical [CNxx, KR86, KRxx, Vis73, Ken80]. Theories [KR86, KRxx, NR97, Cap89]. Theory [AB89, Ozixx, Par14, SHM87, ST89b, UAYS87, De 89, DZ00, HH06, HT91, KK93, KK95a, KK95b, McC94]. thermal [YK90]. Third [Vis73]. Three [BS98, Kra03, Loo72]. Three-loop [BS98]. time [DÜ92a, DÜ92b, Kal82]. times [PH83, SMH98]. Tokyo [WN90]. Tools [KRR88, Par14]. Topology [Ric92]. torsion [Vul02, Vul03]. TR [Mel74]. trace [IKRT91]. traces [IKRT89a]. track [BO80]. trade [Dav82]. trade-off [Dav82]. Transform [Nor80]. Transformation [Edn93, Mag81, Her12]. Transformations [Kaz87, Kaz89]. Transient [SI87]. Transitions [Gro88c]. Translation [WR79]. Translator [Alv00, Mar78]. transport [US91]. treatment [Loo72, Sag88, Sag89]. Triangular [Dul87, Dul89]. Tricks [Kam05]. Tucker [Kam05]. Tutorial [Gri77]. twin [Her12]. Twisted [Kot86]. Two [BOH78]. type [AM90, KKV06].

UK [Fit93b]. Ukraine [Bro93].

Unbounded [Gre87]. Uniform [IK85].

Unit [Zha93]. Univariate

[Mon92, PB90, WD85]. University

[NBC92]. UNIX [HS83]. Untersuchung [GF88]. Unusual [DR86]. upon [Ren92]. Usage [KR86, KRxx, BL91, Lan94]. Use [Gri76, McC84, Nor80, VR87, Ano88, CH85, De 89, Ken80]. User [DJRR81, Gat85c, Gat87, Hea67, Hea68, Hea73b, Hea87, Hea91, SAK^{+}88, vanxx, HM^{+}83]. User-Oriented [Hea68]. Users [Hea74]. Using [Ack85, BBB92, BCBBxx, CB76, DÜ92a, Dul87, Gaw89, GK80a, Gri74, Gro97, Gro05, Hol92, Ioa90a, Kam05, LD87, Mag81, Man93, Mat89, MMN88, Par14, RT85, SAK^{+}88, SKA88, Sto77, SI87, UAYS87, Wan85, dos89, Aok89, BCBB90, Bar92, DÜ92b, Dul89, Her12, IK91, Ioa91a, Kea92, LD89, NM89, Ng89c, Ng89b, Ng89a, ON90, PR02, PBG14, Sag88, Sag89, SMU^{+}89, Ste94, Vul02, Vul03]. Utilisation [Cap84].

vacuum [BS98]. validation [Bam88]. value [Cap86a]. Valued [Asl96]. Variable [Eis90]. Variables [BOH78, Rod84]. VAX [FF82, HS83]. VAX-11 [HS83]. Vaxima [NW83]. Vector [Eas87, Har89b, Eas91, Har89a, Har89a]. Vector33 [Har89b]. Vectors [MN92]. Vers [HS83]. Version [Gat85b, Gat85c, Gat87, Gra94b, Hea91, Eas91, Gra94a, HM^{+}83, Hea87]. Verwendung [GVZ91]. Vibration [OT87, EH88]. Vibration-rotational [OT87]. vibrations [Sag88, Sag89]. vicinity [AZZ91]. Visualization [FKM95, HPR96].

Water [Gro88b]. Waves [Div91, Gro88b]. way [Hie91]. Weak [Gro88c, Gro91b]. Weight [KR86, KRxx]. wire [YK90]. Work [Rod84]. Working [Gon91, Veg91]. Workshop [IEE78, Cal94]. World [HR92]. Writing [Mar78]. Wronskian [Ito88].

Years [Hea05]. York [Jen76]. Yorktown [Jen76].

References

Autin:1989:SEI

[AB89] B. Autin and J. Bengtsson. Symbolic evaluation of integrals occurring in accelerator orbit theory. Journal of Symbolic Computation, 7 (2):183–188 (or 183–187??), February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ackerman:1985:ASE

[Ack85] A. Ackerman. Analysis of a systematic error in field measurements of quadrupole magnets, using the algebraic software package Reduce. Technical Report DESY-M-85-10, ????, ????, 1985. ???? pp.

ACM:1994:IPI

[ACM94] ACM, editor. ISSAC’94. Proceedings of the International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Abi-Ezzi:1983:CSM

[AE83] Salim S. Abi-Ezzi. Clarifications to the symbolic mode in REDUCE. SIGSAM Bulletin, 17(3–4):43–47, August/November 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anderson:1987:IOL

[AGK^{+}87] J. Wayne Anderson, William F.

Galway, Robert R. Kessler, Herbert Melenk, and Winfried Neun. Implementing and optimizing Lisp for the Cray. IEEE Software, 4(4):74–83, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

AlvarezSobreviela:2000:RBO

[Alv00] Luis Alvarez Sobreviela. A Reducebased OpenMath ↔ MathML translator. SIGSAM Bulletin, 34(2):31– 32, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Adamchik:1990:ACI

[AM90] V. S. Adamchik and O. I. Marichev. The algorithm for calculating integrals of hypergeometric type functions and its realization in RE-DUCE system. In Watanabe and Nagata [WN90], pages 212–224. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Anonymous:1988:EPT

[Ano88] Anonymous, editor. ESPRIT ’88: putting the technology to use: proceedings of the 5th Annual ESPRIT Conference, Brussels, November 14–17, 1988. North-Holland, Amsterdam, The Netherlands, 1988. ISBN 0-444-87145-4. LCCN QA75.5 .E84 1988. Two volumes.

Anon:1990:SMD

[Ano90] Anon. Study on a method for direct optical measurement of the nonlinear fracture parameter T · integral. (1st report, numerical simulation aided by computerized symbolic manipulation for the formation process of caustic pattern).

Nippon Kikai Gakkai Ronbunshu, A Hen/Transactions of the Japan Society of Mechanical Engineers, Part A, 56(522):310–316, February 1990. CODEN NKGADA. ISSN 0387-5008.

Aoki:1989:LDA

[Aok89] Genya Aoki. Large deflection analysis by perturbation method using symbolic manipulation system RE-DUCE. Senpaku Gijutsu Kenkyusho Hokoku/Papers of Ship Research Institute, 26(2):21–27, March 1989. CODEN SPGKAP. ISSN 0495-775X.

Aslaksen:1996:MVC

[Asl96] Helmer Aslaksen. Multiple-valued complex functions and computer algebra. SIGSAM Bulletin, 30(2):12– 20, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Antweiler:1989:TSE

[ASW89a] W. Antweiler, A. Strotmann, and V. Winkelmann. A TEX-REDUCE interface. SIGSAM Bulletin, 23 (2):26–33, April 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Antweiler:1989:TRI

[ASW89b] Werner Antweiler, Andreas Strotmann, and Volker Winkelmann. A TEX-REDUCE interface. SIGSAM Bulletin, 23(2):26–33, April 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abdali:1989:EQR

[AW89] S. K. Abdali and D. S. Wiset. Experiments with quadtree representation of matrices. In Gianni [Gia89], pages 96–108. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Amirkhanov:1991:BOV

[AZZ91] I. V. Amirkhanov, E. P. Zhidkov, and I. E. Zhidkova. The betatron oscillations in the vicinity of nonlinear resonance in cyclic accelerator investigation. In Watt [Wat91], pages 452–453. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Bamberger:1988:EVP

[Bam88] L. Bamberger. An error validation package for REDUCE. In Anonymous [Ano88], pages 467–474. ISBN 0-444-87145-4. LCCN QA75.5 .E84 1988. Two volumes.

Barbier:1992:AGB

[Bar92] Christine Barbier. Automatic generation of bending element matrices for finite element method using RE-DUCE. Engineering Computations, 9(4):477–494, August 1992. CO-DEN ENCOEN. ISSN 0264-4401.

Bates:1985:GAC

[Bat85] B. L. Bates. GENTRAN: an automatic code generation facility for REDUCE. SIGSAM Bulletin, 19 (3):24–42, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berkovich:1993:ASS

[BB93] L. M. Berkovich and F. L. Berkovich. Algorithm for solving second order linear ordinary differential equations and its implementation in REDUCE.

Sūrikaisekikenkyūsho Kōkyūroku, 848:100–108, 1993. Theory and applications in computer algebra (Japanese) (Kyoto, 1992).

Barbier:1992:AGM

[BBB92] Christine Barbier, Peter Bettess, and Jacqueline A. Bettess. Automatic generation of mapping functions for infinite elements using RE-DUCE. Journal of Symbolic Computation, 14(5):523–534, November 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bordoni:1981:ARI

[BC81] Luciana Bordoni and Attilio Colagrossi. An application of REDUCE to industrial mechanics. SIGSAM Bulletin, 15(2):8–12, May 1981. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1985:PEE

[BC85] Bruno Buchberger and Bob F. Caviness, editors. Proceedings: EURO-CAL ’85, European Conference on Computer Algebra, Linz, Austria, April 1–3, 1985, volume 203, 204 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNCSD9. ISBN 0-387-15983-5 (v. 1), 0-387-15984-3 (v. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1985. Vol. 2 edited by: Bob F. Caviness. “Jointly organized by the ACM Special Interest Group on Symbolic and Algebraic Manipulation (SIGSAM) and by the Symbolic and Algebraic Manipulation Group in Europe (SAME)”–Vol. 2, pref. Contents: v. 1. Invited lectures — v. 2. Research contributions.

Brackx:1991:CAL

[BC91] F. Brackx and D. Constales. Computer algebra with LISP and REDUCE: an introduction to computer-aided pure mathematics, volume 72 of Mathematics and Its Applications. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1441-7. xii + 264 pp. LCCN QA155.7.E4

                                        B72 1991.  URL http://www.
                                        zentralblatt-math.org/zmath/
                                        en/search/?an=0733.68047. An

introduction to computer-aided pure mathematics.

Barbier:1990:AGS

[BCBB90] Christine Barbier, Philip J. Clark, Peter Bettess, and Jacqueline A. Bettess. Automatic generation of shape functions for finite element analysis using REDUCE. Engineering Computations, 7(4):349–359 (or 349–358??), December 1990. CO-DEN ENCOEN. ISSN 0264-4401.

Barbier:19xx:AGS

[BCBBxx] C. Barbier, P. Clark, P. Bettess, and J. Bettess. Automatic generation of shape functions for finite element analysis using REDUCE. International J. Numer. Methods Engineering, 19xx. submitted.

Brackx:1987:CAR

[BCDS87] Freddy Brackx, Denis Constales, Richard Delanghe, and Herman Serras. Clifford algebra with RE-DUCE. Rend. Circ. Mat. Palermo (2) Suppl., 16:11–19, 1987. Proceedings of the Winter School on Geometry and Physics (Srní, 1987).

Basios:1995:GRP

[BCM^{+}95] V. Basios, N. A. Chekanov, B. L.

Markovski, V. A. Rostovtsev, and S. I. Vinitsky. GITA: A REDUCE program for the normalization of polynomial hamiltonians. Computer Physics Communications, 90(2-3): 355–368, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Brackx:1989:HMD

[BCRS89] F. Brackx, D. Constales, A. Ronveaux, and H. Serras. On the harmonic and monogenic decomposition of polynomials. Journal of Symbolic Computation, 8(3):297– 304, September 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bernardin:1996:RSS

[Ber96] Laurent Bernardin. A review of symbolic solvers. SIGSAM Bulletin, 30(1):9–20, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bennett:1993:CAT

[BGH93] J. P. Bennett, M. Grinfeld, and J. Hubble. Computer algebra techniques in affinity binding equations: the dimer case. Journal of Symbolic Computation, 15(1):79–84 (or 79–83??), January 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boege:1986:SES

[BGK86] W. Boege, R. Gebauer, and H. Kredel. Some examples for solving systems of algebraic equations by calculating Gröbner bases. Journal of Symbolic Computation, 2(1):83– 98, March 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Blum:1993:IOC

[BGS93] W. Blum, V. Ganzha, and W. Strampp. An introduction to ODE’s by CAS. In G. Jacob, N. E. Oussous, and S. Steinberg, editors, Proceedings of the 1993 International IMACS Symposium on Symbolic Computation, pages 110–119. IMACS, Laboratoire d’Informatique Fondamentale de Lille, France, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1993.

Borst:1994:GRP

[BGV94] W. N. Borst, V. V. Goldman, and J. A. Van Hulzen. GENTRAN 90: a REDUCE package for the generation of Fortran 90 code. In ACM [ACM94], pages 45–51. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Bogdanova:1988:RPE

[BH88] N. Bogdanova and H. Hogreve. A REDUCE package for exact Coulomb interaction matrix elements. Computer Physics Communications, 48(2):319–326, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Bradford:1986:ERD

[BHPS86] R. J. Bradford, A. C. Hearn, J. A.

Padget, and E. Schrüfer. Enlarging the REDUCE domain of computation. In Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation, pages 100–106. ????, ????, July 21–23, 1986.

Bilge:1992:RPI

[Bil92] Ayşe Hümeyra Bilge. A RE-DUCE program for the integration of differential polynomials. Computer Physics Communications, 71 (3):263–268, September 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Broughan:1991:SSC

[BKR^{+}91] K. A. Broughan, G. Keady, T. D.

Robb, M. G. Richardson, and M. C. Dewar. Some symbolic computing links to the NAG numeric library. SIGSAM Bulletin, 25(3):28– 37, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

BelKov:1991:RUC

[BL91] Alexander A. Bel’Kov and Alexander V. Lanyov. REDUCE usage for calculation of low-energy process amplitudes in chiral QCD model. In Watt [Wat91], pages 454–455. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Bajla:1980:ARA

[BO80] I. Bajla and G. A. Ososkov. Application of REDUCE-2 algebraic manipulation system in calibration problems of track chamber picture processing devices. Computer Physics Communications, 20(1):81– 83, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0010465580901125.

Bocko:1992:ERB

[Boc92] J. Bocko. EQSHELL — a REDUCE-based program for generation of equations of equilibrium for shells. Computer Physics Communications, 69(1):215–222, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Bajla:1978:OPP

[BOH78] I. Bajla, G. A. Ososkov, and A. C. Hearn. The orthogonalization program of polynomials in two variables in REDUCE-2 language. Report P10-11944, J.I.N.R., Dubna, USSR, 1978.

Boyd:1993:CLS

[Boy93] John P. Boyd. Chebyshev and Legendre spectral methods in algebraic manipulation languages. Journal of Symbolic Computation, 16(4): 377–399, October 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bradford:1990:PBA

[Bra90] Russell Bradford. Parallelization of the Buchberger algorithm. In ISSAC ’90 Proceedings of International Symposium on Symbolic and Algebraic Computation (Aug 20–24 1990: Tokyo, Jpn), pages 296–?? ACM, New York, NY, USA, 1990. ISBN 0-201-54892-5. LCCN ????

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. IS-SAC’93: proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6– 8, 1993, Kiev, Ukraine. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Brunswick:1993:LAR

[Bru93] B. D. Brunswick. LILK: static analysis of REDUCE code. In Fitch [Fit93b], pages 35–43. CO-DEN LNCSD9. ISBN 3-540-57272-4 (Berlin), 0-387-57272-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1992.

Baikov:1998:TLV

[BS98] P. A. Baikov and M. Steinhauser. Three-loop vacuum integrals in FORM and RE-DUCE. Computer Physics Communications, 115(2–3):161–169, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0010465598001258.

Calmet:1972:RAC

[Cal72] Jacques Calmet. A REDUCE approach to the calculation of Feynman diagrams. Computer Physics Communications, 4(2):199– 204, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Calmet:1994:RWC

[Cal94] J. Calmet, editor. Rhine workshop on computer algebra: March 22– 24, 1994, Karlsruhe, Germany. Universität Karlsruhe, Karlsruhe, Germany, 1994. ISBN ???? LCCN ????

Caprasse:1984:DUD

[Cap84] H. Caprasse. Description et utilisation d’une extension du programme REDUCE. Technical report, Physique Théorique et Mathématique, Université de Liège, Liège, Belgium, October 1984.

Caprasse:1986:CSP

[Cap86a] H. Caprasse. A complete simplification package for the absolute value function in REDUCE. SIGSAM Bulletin, 20(1 and 2):18–21, February/May 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Implementation for REDUCE 3.2 of the function ABS.

Caprasse:1986:DEM

[Cap86b] H. Caprasse. Description of an extension of the matrix package of “Reduce”. SIGSAM Bulletin, 20 (4):7–10, December 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caprasse:1989:TCM

[Cap89] H. Caprasse. Les théories des Champs dans le monde de RE-DUCE (in French). CALSYF, 1989.

Caprasse:1990:RGF

[Cap90] H. Caprasse. Renormalization group, function iterations and computer algebra. Journal of Symbolic Computation, 9(1):61–72, January 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cohen:1976:PCS

[CB76] I. Cohen and F. Bark. Perturbation calculations for the spin up problem using REDUCE. Technical Report TRITA-MEK-76-03, The Royal Institute of Technology, Department of Mechanics, ????, 1976.

Chaffy-Camus:1988:ARA

[CC88] C. Chaffy-Camus. An application of REDUCE to the approximation of f(x, y). In Proc. of ISSAC ’88, volume 358, pages 73–84. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988.

Chaffy-Camus:1989:ARA

[CC89] C. Chaffy-Camus. An application of REDUCE to the approximation of f(x, y). In Gianni [Gia89], pages 73–84. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Caprasse:1985:NUO

[CH85] H. Caprasse and M. Hans. A new use of operators in the algebraic mode of REDUCE. SIGSAM Bulletin, 19(3):46–52, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Calmet:1996:DIS

[CL96] Jacques Calmet and Carla Limongelli, editors. Design and implementation of symbolic computation systems: International Symposium, DISCO ’96, Karlsruhe, Germany, September 18–20, 1996: proceedings, volume 1128 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996.

Cejchan:19xx:ARA

[CNxx] A. Cejchan and J. Nadrchal. Application of REDUCE-2 and analytic integration program in the theoretical solid state physics. Technical report, Institute of Physics, CSAV, Prague, Czechoslovakia, 19xx.

Dautcourt:1979:ARA

[Dau79] G. Dautcourt. Application of RE-DUCE to algebraic computations in general relativity and astrophysics. In Proc. of the Workshop in Symbolic Computation, Dubna, USSR. ????, ????, September 1979.

Davenport:1982:FRT

[Dav82] James H. Davenport. Fast RE-DUCE: The trade-off between efficiency and generality. SIGSAM Bulletin, 16(1):8–11, February 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1988:CAA

[Dav88] J. H. Davenport. Computer algebra applied to itself. Journal of Symbolic Computation, 6(1): 127–132, August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Davenport:1989:EEC

[Dav89] James Harold Davenport, editor.

EUROCAL ’87: European Conference on Computer Algebra, Leipzig, GDR, June 2–5, 1987: proceedings, volume 378 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

DeVos:1989:URS

[De 89] Alexis De Vos. The use of Reduce in solar energy conversion theory. Reports of the CAGe Project 4, State University of Gent, CAGe Computer Algebra Group, Gent, Belgium, August 1989.

Dewar:1989:IIS

[Dew89] M. C. Dewar. IRENA: an integrated symbolic and numerical computation environment. In Gonnet [Gon89], pages 171–179. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Diver:1991:MWC

[Div91] D. A. Diver. Modelling waves with computer algebra. Journal of Symbolic Computation, 11(3):275– 290 (or 275–289??), March 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dautcourt:1980:PPR

[DJ80] G. Dautcourt and K. P. Jann. A program package in REDUCE 2 for algebraic computations in general relativity. Technical report, Zentralinstitut für Astrophysik der Akademie der Wissenschaften, ????, 1980.

Dudley:1989:CAD

[DJ89] M. L. Dudley and R. W. James. Computer-aided derivation of spherical harmonic spectral equations in astro-geophysics. Journal of Symbolic Computation, 8(4):423–427, October 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dautcourt:1981:UGR

[DJRR81] G. Dautcourt, K. P. Jann, E. Riemer, and M. Riemer. User’s guide to REDUCE subroutines for algebraic computations in general relativity. Astron. Nachr., 302:1– 13, 1981. CODEN ASNAAN. ISSN 0004-6337.

Darbaidze:1988:SCR

[DMR88] Ya. Z. Darbaidze, Z. V. Merebashvili, and V. A. Rostovtsev. Some computer realizations of the REDUCE-3 calculations for exclusive processes. Preprint P2-88-769, J.I.N.R., Dubna, USSR, 1988.

Darbaidze:1990:SCR

[DMR90] Ya. Z. Darbaidze, Z. V. Merebashvili, and V. A. Rostovtsev. Some computer realizations of the REDUCE-3 calculations for exclusive processes. Fortschritte der Physik = Progress of Physics, 38 (9):717–731, ???? 1990. CODEN FPYKA6. ISSN 0015-8208.

DeGroot:1993:PCP

[DN93] Robert A. De Groot and Jaroslav Nadrchal, editors. Physics computing ’92: proceedings of the 4th international conference, Prague, Czechoslovakia, August 24–28, 1992, Physics Computing — International Conference. World Scientific Publishing, Singapore, 1993. ISBN 981-02-1245-3. LCCN QC19.2.I53 1992.

dosSantos:1988:IAS

[dos88] Renato P. dos Santos. Introdução ao sistema REDUCE de Cálculo algébrico. Technical Report CBPF-NT-001/88, CBPF, Rio de Janeiro, Brazil, 1988.

dosSantos:1989:URS

[dos89] Renato P. dos Santos. Using RE-DUCE in supersymmetry. Journal of Symbolic Computation, 7(5):523– 525, May 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Demichev:1985:RPC

[DR85] A. P. Demichev and A. Ya. Rodinov. A REDUCE program for the calculation of geometrical characteristics of compactified multidimensional Riemannian space. Computer Physics Communications, 38: 441–448, 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Duncan:1986:RUM

[DR86] Anthony Duncan and Ralph Roskies.

Representations of unusual mathematical structures in scientific applications of symbolic computation.

Journal of Symbolic Computation, 2(2):201–206 (or 201–207??), June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dolzmann:1997:RCA

[DS97] Andreas Dolzmann and Thomas Sturm. REDLOG: Computer algebra meets computer logic. SIGSAM Bulletin, 31(2):2–9, June 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dereli:1992:SSA

[DÜ92a] T. Dereli and G. Üçoluk. A study of stationary, axially symmetric spacetime geometries satisfying modified double duality equations using the exterior calculus package X^{T R} for REDUCE. Journal of Computational Physics, 99(1):169– 175, March 1992. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0021999192902846.

Derili:1992:SSA

[DÜ92b] T. Derili and G. Üçoluk. A study of stationary, axially symmetric space-time geometries satisfying modified double duality equations using the exterior calculus package X^{T} R for Reduce. Journal of Computational Physics, 98 (2):349, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/002199919290163S.

Dulyan:1987:CQT

[Dul87] L. S. Dulyan. The calculation of QCD triangular Feynman graphs in the external gluonic field using REDUCE-2 system. In Proc. EU-ROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 172–173. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Dulyan:1989:CQT

[Dul89] L. S. Dulyan. The calculation of QCD triangular Feynman graphs in the external gluonic field using REDUCE-2 system. In Davenport [Dav89], pages 172–173. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

Dyer:1994:ASC

[Dye94] C. C. Dyer. An application of symbolic computation in the physical sciences. In ACM [ACM94], pages 181–186. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Denisova:2000:ACA

[DZ00] I. P. Denisova and A. A. Zubrilo.

Application of the computer algebra “REDUCE” for integrating equations of gravitation theory by the Newman–Penrose method. Mat. Model., 12(2):59–67, 2000. ISSN 0234-0879.

Eastwood:1987:OAR

[Eas87] James W. Eastwood. ORTHOVEC:

A REDUCE program for 3-D vector analysis in orthogonal curvilinear coordinates. Computer Physics Communications, 47(1):139–147, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Eastwood:1991:OVR

[Eas91] James W. Eastwood. ORTHOVEC: version 2 of the REDUCE program for 3-D vector analysis in orthogonal curvilinear coordinates. Computer Physics Communications, 64 (1):121–122, April 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Elishakoff:1987:ASA

[EC87] Isaac Elishakoff and Brian Couch. Application of symbolic algebra to the instability of a nonconservative system. Journal of Symbolic Computation, 4(3):391–396, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Edelen:1981:PCI

[Ede81] Dominic G. B. Edelen. Programs for calculation of isovector fields in the REDUCE-2 environment. Technical Report TBD, Center for the Application of Mathematics, Lehigh University, ????, August 1981.

Edneral:1993:CGN

[Edn93] V. F. Edneral. Computer generation of normalizing transformation for systems of nonlinear ODE. In Bronstein [Bro93], pages 14–19. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Eliseev:1985:RPD

[EFK85] V. P. Eliseev, R. N. Fedorova, and V. V. Kornyak. A REDUCE program for determining point and contact Lie symmetries of differential equations. Computer Physics Communications, 36(4):383–389, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Elishakoff:1988:NRV

[EH88] Isaac Elishakoff and Charles D. Hettema. Nonstationary random vibration with REDUCE. In P. D. (Pol D.) Spanos, editor, Probabilistic Methods in Civil Engineering (May 25–27 1988: Blacksburg, VA, USA), pages 529–532. ASCE, New York, NY, USA, 1988. ISBN 0-87262-659-8. LCCN TA340 .P7351 1988.

Eisenberger:1990:ASA

[Eis90] Moshe Eisenberger. Application of symbolic algebra to the analysis of plates on variable elastic foundation. Journal of Symbolic Computation, 9(2):207–214 (or 207–213??), February 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Edneral:1989:LAC

[EKR89] Viktor F. Edneral, Aleksandr P.

Kryukov, and Anatolii Ia. Rodinov. The language of the analytic computer program REDUCE. Moscow, Izd-vo, Moskovskogo unta, Moscow, USSR, 1989. ISBN ???? ???? pp. LCCN ????

Elishakoff:1990:CSA

[EP90] Isaac Elishakoff and Baruch Pletner. Computerized symbolic algebraic evaluation of buckling loads by Snitko’s method. American Society of Mechanical Engineers, Pressure Vessels and Piping Division (Publication) PVP, 205:175–187, 1990. CODEN AMPPD5. ISSN 0277-027X.

Elishakoff:1991:ABC

[EP91] Isaac Elishakoff and Baruch Pletner. Analysis of buckling by computer algebra. Computer Methods in Applied Mechanics and Engineering, 88(3):299–309, July 1991. CO-DEN CMMECC. ISSN 0045-7825, 0374-2830.

Eldeib:1989:ASM

[ET89] H. K. Eldeib and S. Tsai. Applications of symbolic manipulation in control system analysis and design. In IEEE, editor, IEEE International Symposium on Intelligent Control, 1988. Proceedings, 24–26 August 1988, Arlington, Virginia, pages 269–274. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-2012-9. LCCN TJ212.2 .I2 1988.

Flatau:1986:SAA

[FBC86] Piotr J. Flatau, John P. Boyd, and William R. Cotton. Symbolic algebra in applied mathematics and geophysical fluid dynamics — RE-DUCE examples. Technical report, Dept. of Atmospheric and Oceanic Science, University of Michigan, and Dept. of Atmospheric Science, Colorado State University, Ann Arbor, MI, USA and Ft. Collins, CO, USA, 1986.

Fogelholm:1982:SLV

[FF82] Rabbe Fogelholm and Inge B. Frick.

Standard LISP for the VAX: a provisional implementation. SIG-SAM Bulletin, 16(4):10–12, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Freire:1989:ASC

[FGPF89] E. Freire, E. Gamero, E. Ponce, and L. G. Franquelo. An algorithm for symbolic computation of center manifolds. In Gianni [Gia89], pages 218–230. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Fitch:1973:PRM

[Fit73] John Fitch. Problems #3 and #4 in REDUCE and MACSYMA. SIG-SAM Bulletin, ??(??):10–11, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1983:IRM

[Fit83] J. P. Fitch. Implementing RE-DUCE on a microprocessor. In Proc. EUROCAL 1983, Lecture Notes in Computer Science, volume 162, pages 128–136. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983.

Fitch:1985:SAP

[Fit85] John P. Fitch. Solving algebraic problems with REDUCE. Journal of Symbolic Computation, 1(2):211– 228 (or 211–227??), June 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fitch:1989:CRB

[Fit89] J. Fitch. Can REDUCE be run in parallel? In Gonnet [Gon89], pages 155–162. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Fitch:1990:DSR

[Fit90] John Fitch. A delivery system for REDUCE. In Watanabe and Nagata [WN90], pages 76–81. ISBN 0-201-54892-5. LCCN ????

Fitch:1993:RMC

[Fit93a] J. P. Fitch. REDUCE meets CA-MAL. In Fitch [Fit93b], pages 104– 15. CODEN LNCSD9. ISBN 3-540-57272-4 (Berlin), 0-387-57272-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1992.

Fitch:1993:DIS

[Fit93b] John Fitch, editor. Design and implementation of symbolic computation systems: International Symposium, DISCO ’92, Bath, UK, April 13–15, 1992: proceedings, volume 721 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CO-DEN LNCSD9. ISBN 3-540-57272-4 (Berlin), 0-387-57272-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1992.

Feldmar:1986:RPM

[FK86] E. Feldmar and K. S. Kölbig. Reduce procedures for the manipulation of generalized power series. Computer Physics Communications, 39(2):267–284, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Fushchich:1989:CAA

[FK89] W. I. Fushchich and V. V. Kornyak.

Computer algebra application for determining Lie and Lie–Bäcklund symmetries of differential equations. Journal of Symbolic Computation, 7 (6):611–619, June 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fournier:1995:VMS

[FKM95] Robert Fournier, Norbert Kajler, and Bernard Mourrain. Visualization of mathematical surfaces: the IZIC server approach. Journal of Symbolic Computation, 19(1/2/3): 159–174 (or 159–173??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Gates:1976:NCG

[Gat76] B. L. Gates. A numerical code generation facility for REDUCE. In Jenks [Jen76], pages 94–99. LCCN QA155.7.E4 .A15 1976. US$20.00.

Gates:1985:GAC

[Gat85a] Barbara L. Gates. GENTRAN: an automatic code generation facility for REDUCE. SIGSAM Bulletin, 19 (3):24–42, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gates:1985:GDI

[Gat85b] Barbara L. Gates. Gentran design and implementation, REDUCE version. Memorandum Memorandum INF-85-12, Twente University of Technology, Department of Computer Science, Enschede, The Netherlands, August 1985.

Gates:1985:GUM

[Gat85c] Barbara L. Gates. Gentran user’s manual — REDUCE version. Memorandum INF-85-11, Twente University of Technology, Department of Computer Science, Enschede, The Netherlands, June 1985.

Gates:1986:NCG

[Gat86] Barbara L. Gates. A numerical code generation facility for REDUCE. In Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation, pages 94–99. ????, ????, July 21–23, 1986.

Gates:1987:GUM

[Gat87] B. L. Gates. GENTRAN User’s Manual, REDUCE Version. The RAND Corporation, Santa Monica, CA, USA, 1987.

Gawthrop:1989:BGA

[Gaw89] Peter Gawthrop. Bond graph analysis using Prolog and Reduce. Control Group Report (89/7), Engineering Design Research Centre, ????, September 1989.

Ganzha:1988:KUD

[GF88] Victor G. Ganzha and Reinhard Foessmeier. Konstruktion und Untersuchung von Differenzenoperationen mit dem Reduce-system. Technical report, ????, München, Germany, 1988.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4–8, 1988: proceedings, volume 358 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Griss:1978:RMA

[GK78a] M. L. Griss and R. R. Kessler. RE-DUCE/ 1700: a micro-coded algebra system. In IEEE [IEE78], pages 130–138. ISSN 0194-1895. LCCN QA76.6 .W69a. Catalog number 78 CH1411-8.

Griss:1978:RAM

[GK78b] Martin L. Griss and Robert R.

Kessler. REDUCE/1700: a microcoded algebra system. In Proc. Micro, IEEE, volume 11, pages 130– 138. ????, ????, 1978.

Griss:1979:MIS

[GK79] Martin L. Griss and Robert R.

Kessler. A micro-programmed implementation of Standard LISP and REDUCE on the Burroughs B1700/B1800 computer. Report, University of Utah, Salt Lake City, UT, USA, February 1979.

Grimm:1980:URP

[GK80a] R. Grimm and H. Kühnelt. Using REDUCE in problems of supersymmetry and supergravity. Computer Physics Communications, 20: 77, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Griss:1980:RSL

[GK80b] M. L. Griss and R. R. Kessler. RE-DUCE/1800: Standard LISP and REDUCE on the Burroughs B1700/ B1800 computer. Report, Computer Science Department, University of Utah, Salt Lake City, UT, USA, June 1980.

Gragert:1991:GDG

[GK91] P. K. H. Gragert and P. H. M. Kersten. Graded differential geometry in REDUCE: Supersymmetry structures of the modified KdV equation. Acta Applicandae Mathematicae, 24 (3):211–232, September 1991. CO-DEN AAMADV. ISSN 0167-8019.

Gao:2002:CCQ

[GKI02a] Min Gao, Yasuyuki Kato, and Masaaki Ito. Calculation of the conserved quantity for rational difference equations through REDUCE. Sūrikaisekikenkyūsho Kōkyūroku, 1295:56–61, 2002. Computer algebra—algorithms, implementations and applications (Kyoto, 2001).

Gao:2002:RPF

[GKI02b] Min Gao, Yasuyuki Kato, and Masaaki Ito. A REDUCE package for finding conserved densities of systems of nonlinear differencedifference equations. Computer Physics Communications, 148(2): 242–255, 2002. CODEN CPHCBZ. ISSN 0010-4655.

Golovko:2008:ICH

[GKKV08] V. Golovko, P. Kersten, I. Krasil′shchik, and A. Verbovetsky. On integrability of the Camassa–Holm equation and its invariants: a geometrical approach. Acta Applicandae Mathematicae, 101(1–3):59–83, 2008. CO-DEN AAMADV. ISSN 0167-8019.

Grabmeier:2003:CAH

[GKW03] Johannes Grabmeier, Erich Kaltofen, and Volker Weispfenning, editors. Computer algebra handbook: foundations, applications, systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-65466-6. xx + 637 pp. LCCN QA155.7.E4 C64954 2003.

URL http://www.springer.com/
sgw/cda/frontpage/0,11855,1-
102-22-1477871-0,00.html. In-

cludes CD-ROM.

Gerdt:1991:LSC

[GKZ91] V. P. Gerdt, N. V. Khutornoy, and A. Yu. Zharkov. Lie-Backlund symmetries of coupled nonlinear Schrodinger equations. In Watt [Wat91], pages 313–314. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Ganzha:1988:ARC

[GL88] Victor G. Ganzha and Richard Liska. Application of the Reduce computer algebra system to stability analysis of difference schemes. Technical report, ????, München, Germany, 1988.

Ganzha:1989:ARC

[GL89] Victor G. Ganzha and Richard Liska. Application of the REDUCE computer algebra system to stability analysis of difference schemes. In E. Kaltofen and S. M. Watt, editors, Proc. Computers and Mathematics ’89, pages 119–129. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989.

Gebauer:1988:IBA

[GM88] Rüdiger Gebauer and H. Michael Möller. On an installation of Buchberger’s algorithm. Journal of Symbolic Computation, 6(2–3):275–286, October/December 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Ganzha:1990:ARS

[GMS90] V. G. Ganzha, S. V. Meleshko, and V. P. Shelest. Application of REDUCE system for analyzing consistency of systems of PDE’s. In Watanabe and Nagata [WN90], pages 301–?? ISBN 0-201-54892-5. LCCN ????

Ganzha:1993:ARS

[GMS93] V. G. Ganzha, S. V. Meleshko, and V. P. Shapeev. Application of the REDUCE system to the analysis of the consistency of systems of differential equations. Model. Mekh., 7 (4):26–35 (1994), 1993. ISSN 0235-

Gonnet:1989:PAI

[Gon89] Gaston H. Gonnet, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89 / July 17–19, 1989, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

GonzalezVega:1991:WRA

[Gon91] L. Gonzalez Vega. Working with real algebraic plane curves in RE-DUCE: the GCUR package. In Watt [Wat91], pages 397–402. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Grabe:1993:LP

[Grä93] Hans-Gert Gräbe. On lucky primes. Journal of Symbolic Computation, 15(2):199–210 (or 199– 209??), February 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Grabe:1994:SDC

[Gra94a] H.-G. Grabe. A short description of CALI — a REDUCE package for commutative algebra (version 2.1). In Calmet [Cal94], pages 182–193. ISBN ???? LCCN ????

Graebe:1994:SDC

[Gra94b] H.-G. Graebe. A short description of CALI — a REDUCE package for commutative algebra (version 2.1). In Calmet [Cal94], pages 182–193. ISBN ???? LCCN ????

Grabe:1995:ALA

[Grä95] Hans-Gert Gräbe. Algorithms in local algebra. Journal of Symbolic Computation, 19(6):545–558 (or 545–557??), June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Greenberg:1987:CAA

[Gre87] Harvey J. Greenberg. Computerassisted analysis for diagnosing infeasible or unbounded linear programs. Mathematical Programming Study, 31(??):79–97, September 1987. CODEN MPSTDF. ISSN 0303-3929.

Griss:1974:ASL

[Gri74] M. L. Griss. The algebraic solution of large sparse systems of linear equations using REDUCE 2. In Proceedings of ACM 74, pages 105–111. ????, ????, 1974.

Griss:1975:RSC

[Gri75] Martin L. Griss. The REDUCE system for computer algebra. In Proc. ACM 75, pages 4–5. ????, ????, 1975.

Griss:1976:DUD

[Gri76] Martin L. Griss. The definition and use of data-structures in REDUCE. In Proc. of the 1976 ACM Symposium on Symbolic and Algebraic Computation, pages 53–59. ACM Press, New York, NY 10036, USA, 1976.

Griss:1977:RST

[Gri77] Martin L. Griss. A REDUCE symbolic-numeric tutorial. Utah Symbolic Computation Group Operating Note UCP-32, Department of Computer Science, University of Utah, Salt Lake City, UT, USA, October 1977.

Griss:1981:RSS

[Gri81] Martin L. Griss. REDUCE.SYSLISP.

SOURCES, BIGNUM package. Report, Univ. of Utah, CS Dept., Salt Lake City, UT, USA, August 1981. Recoded from BIGNUM.MOD, Dec. 1977.

Grozin:1983:COD

[Gro83] A. G. Grozin. Calculation of oneloop diagrams of 1 → 2 decays with REDUCE. In Proc. Int. Conf. on Computer Algebra in Theoretical Physics, Dubna, pages 226–231. ????, ????, 1983.

Grozin:1988:SPPa

[Gro88a] A. G. Grozin. Solving physical problems with REDUCE. 1. REDUCE language 2. Classical nonlinear oscillator. Preprint 88-115, Institute of Nuclear Physics, 630090, Novosibirsk, USSR, 1988.

Grozin:1988:SPPb

[Gro88b] A. G. Grozin. Solving physical problems with REDUCE. 3. Nonlinear water waves 4. Calculation of the curvature tensor 5. Angular momentum addition. Preprint 88-136, Institute of Nuclear Physics, 630090, Novosibirsk, USSR, 1988.

Grozin:1988:SPPc

[Gro88c] A. G. Grozin. Solving physical problems with REDUCE. 6. Quantum nonlinear oscillator 7. Rotator in a weak field 8. Radiative transitions in charmonium. Preprint 88-140, Institute of Nuclear Physics, 630090, Novosibirsk, USSR, 1988.

Grozin:1990:REPa

[Gro90a] A. G. Grozin. REDUCE in elementary particle physics. introduction. Technical Report INP 90-42, Institute of Nuclear Physics, Novosibirsk, USSR, 1990.

Grozin:1990:REPc

[Gro90b] A. G. Grozin. REDUCE in elementary particle physics. quantum chromodynamics. Technical Report INP 90-62, Institute of Nuclear Physics, Novosibirsk, USSR, 1990.

Grozin:1990:REPb

[Gro90c] A. G. Grozin. REDUCE in elementary particle physics. quantum electrodynamics. Technical Report INP 90-71, Institute of Nuclear Physics, Novosibirsk, USSR, 1990.

Grozin:1991:REPa

[Gro91a] A. G. Grozin. REDUCE in elementary particle physics. radiative corrections. Technical Report INP 91-46, Institute of Nuclear Physics, Novosibirsk, USSR, 1991.

Grozin:1991:REPb

[Gro91b] A. G. Grozin. REDUCE in elementary particle physics. weak interactions. Technical Report INP 91-56, Institute of Nuclear Physics, Novosibirsk, USSR, 1991.

Grozin:1997:URH

[Gro97] A. G. Grozin. Using REDUCE in high energy physics. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-01952-4, 0-521-56002-0 (hardcover). xiv + 384 pp. LCCN QC793.3.H5 G76 1997. URL http:

//www.inp.nsk.su/persons/A_G_

Grozin/book/. English translation of an original Russian edition.

Grozin:2005:URH

[Gro05] A. G. Grozin. Using REDUCE in high energy physics. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-01952-4 (paperback). xiii + 384 pp. LCCN QC793.3.H5

G76 2005. URL http://www.
loc.gov/catdir/enhancements/
fy0632/2006272924-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0632/2006272924-t.html.

Ganzha:1990:LAS

[GS90] Victor G. Ganzha and Michail Yu.

Shashkov. Local approximation study of difference operations by means of REDUCE system. In Watanabe and Nagata [WN90], pages 185–192. ISBN 0-201-54892-

5. LCCN ????

Gerdt:1985:CAA

[GSZ85] V. P. Gerdt, A. B. Shvachka, and A. Yu. Zharkov. Computer algebra application for classification of integrable non-linear evolution equations. Journal of Symbolic Computation, 1(1):101–107, March 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gerdt:1991:RPS

[GT91] V. P. Gerdt and P. Tiller. A Reduce program for symbolic computation of Puiseux expansions. Preprint E5-91-401, Joint Institute for Nuclear Research, Dubna, Russia, 1991. 9 pp. With a Russian summary.

Gruber:1994:PJE

[GT94] Ralf Gruber and Marco Tomassini, editors. Proceedings of the 6th Joint EPS-APS International Conference on Physics Computing: Physics Computing ’94, Palazzo dei Congressi, Lugano, Switzerland, 22– 26 August 1994. European Physical Society, Geneva, Switzerland, 1994. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Ganzha:1992:RBC

[GV92] V. G. Ganzha and E. V. Vorozhtsov. Resultant based code generation for the stability analysis of difference schemes with the aid of the RE-DUCE system. In Houstis and Rice [HR92], pages 185–192. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Ganzha:1991:SDM

[GVZ91] Victor G. Ganzha, Evgenii V. Vorozhtsov, and Christoph Zenger. Stabilitätsuntersuchung von Differenzenverfahren mit Hilfe der Resultantenalgebra unter Verwendung des Reduce-systems. Technical Report TUM-I9120, Institute for Computer Science, Technical University Munich, München, Germany, 1991.

Gerdt:1983:RPS

[GZ83] V. P. Gerdt and A. Yu. Zharkov. REDUCE — a package for solution of ordinary differential equations. In Dubna 1983. ????, ????, 1983.

Gerdt:1990:CCI

[GZ90a] V. P. Gerdt and A. Yu. Zharkov.

Computer classification of integrable coupled KdV-like systems. Journal of Symbolic Computation, 10(2):203–208 (or 203–207??), August 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gerdt:1990:CGN

[GZ90b] Vladimir P. Gerdt and Alexey Yu.

Zharkov. Computer generation of necessary integrability conditions for polynomial-nonlinear evolution systems. In Watanabe and Nagata [WN90], pages 250–254. ISBN 0-201-54892-5. LCCN ????

Harrington:1977:RSP

[Har77a] S. J. Harrington. REDUCE solution to problem #8. SIGSAM Bulletin, 11–12(4–1):7–8, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Harrington:1977:SLE

[Har77b] Steven J. Harrington. A symbolic limit evaluation program in RE-DUCE. A program for the automatic evaluation of algebraic limits, implemented in MODE-REDUCE, is described. The program incorporates many of the techniques previously employed, including the topdown recursive evaluation, power series expansion, and L’Hopital’s rule. It also introduces the concept of a special algebraic form for limits., 1977.

Harrington:1979:NSI

[Har79a] Steven J. Harrington. A new symbolic integration system in RE-DUCE. The Computer Journal, 22 (2):127–131, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Harrington:1979:SLE

[Har79b] Steven J. Harrington. A symbolic limit evaluation program in RE-DUCE. SIGSAM Bulletin, 13(1): 27–31, February 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Haraldsson:1980:RSF

[Har80] A. Haraldsson. REDUCE — sprak foer formelbehandling. Report 1980-02-20, Linköping Universität, Tekniska Högskolan, Datalogi, Linköping, February 1980.

Harper:1989:VRP

[Har89a] D. Harper. VECTOR 33: a REDUCE program for vector algebra and calculus in orthogonal curvilinear coordinates. Computer Physics Communications, 54(2-3): 295–305, June–July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Harper:1989:VAR

[Har89b] David Harper. Vector33: A RE-DUCE program for vector algebra and calculus in orthogonal curvilinear coordinates. Computer Physics Communications, 54(2 and 3):295– 305, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Hartley:1996:CSI

[Har96] D. Hartley. Causal structure and integrability in moving frames with Reduce. In Hehl et al. [HPR96], pages 270–284. ISBN 3-540-60361-1. LCCN QC173.55 .R445 1996.

Hearn:1967:RUM

[Hea67] A. C. Hearn. REDUCE User’s Manual. Stanford University, Department of Computer Science, Stanford, CA, USA, February 1967. ???? pp. Memo AI-50.

Hearn:1968:RAU

[Hea68] Anthony C. Hearn. REDUCE: a user-oriented interactive system for algebraic simplification. In Klerer and Reinfelds [KR68], page ?? LCCN QA76.5.A1 A2 1967. ACM Symposium on Interactive Systems for Experimental Mathematics. Washington, DC, USA.

Hearn:1969:R

[Hea69] Anthony C. Hearn. REDUCE. Report DECUS No. 10-21, Univ. of Utah, Dept. of Physics, Salt Lake City, UT, USA, October 1969.

Hearn:1971:RAS

[Hea71] Anthony C. Hearn. REDUCE 2 — a system and language for algebraic manipulation. In S. R. Petrick, editor, Proceedings of the Second ACM Symposium on Symbolic and Algebraic Manipulation, pages 128–133. ACM Press, New York, NY 10036, USA, 1971.

Hearn:1972:R

[Hea72a] A. C. Hearn. REDUCE. SIG-SAM Bulletin, ??(??):14–15, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hearn:1972:RSP

[Hea72b] Anthony C. Hearn. A REDUCE solution of problem #2 — the Y(2n) functions. SIGSAM Bulletin, 14, 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hearn:1973:RPC

[Hea73a] A. C. Hearn. The REDUCE program for computer algebra. In Visconti [Vis73], pages AV/1–AV/19. ISBN ???? LCCN ???? Two volumes.

Hearn:1973:RUM

[Hea73b] Anthony C. Hearn. REDUCE 2 user’s manual. Utah Symbolic Computation Group Report UCP-19, Department of Computer Science, University of Utah, Salt Lake City, UT, USA, March 1973.

Hearn:1974:RUM

[Hea74] A. C. Hearn. REDUCE 2, Users manual. Salt Lake City, UT, USA, 1974.

Hearn:1976:NRM

[Hea76] Anthony C. Hearn. A new RE-DUCE model for algebraic simplification. In Proceedings of the 1976 ACM Symposium on Symbolic and Algebraic Computation, pages 46– 52. ACM Press, New York, NY 10036, USA, 1976.

Hearn:1982:RCS

[Hea82] Anthony C. Hearn. Reduce — a case study in algebra system development. Lecture Notes in Computer Science, pages 263–272, 1982. CO-DEN LNCSD9. ISBN 3-540-11607-9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hearn:1987:RUM

[Hea87] Anthony C. Hearn. REDUCE user’s manual, Version 3.3. Report CP 78, The RAND Corporation, Santa Monica, CA, USA, July 1987.

Hearn:1991:RUM

[Hea91] A. C. Hearn. REDUCE User’s Manual, Version 3.4. The Rand Corporation, Santa Monica (CA), 1991.

Hearn:2005:RFF

[Hea05] Anthony C. Hearn. REDUCE: The first forty years. In Andreas Dolzmann, Andreas Seidl, and Thomas Sturm, editors, Algorithmic Algebra and Logic: Proceedings of the A3L 2005, April 3–6, Passau, Germany, Conference in Honor of the 60th Birthday of Volker Weispfenning, pages 19–24. Herstellung und Verlag: Books on Demand GmhH, Norderstedt, Germany, 2005. ISBN 3-8334-2669-1. LCCN A155.7.E4

A39 2005. URL http://reduce-
algebra.com/reduce40.pdf.

Hearn:2009:RFS

[Hea09] Anthony C. Hearn. REDUCE is free software as of January 2009. ACM Communications in Computer Algebra, 43(1):15–16, March 2009. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hermanto:2012:ETP

[Her12] Arief Hermanto. The explanation of the twin paradox using Poincaré transformation and computer algebra system REDUCE. AIP Conference Proceedings, 1454 (1):186–188, 2012. URL http:

//scitation.aip.org/content/
aip/proceeding/aipcp/10.1063/
1.4730717.

Hanzon:2006:CAS

[HH06] Bernard Hanzon and Michiel Hazewinkel, editors. Constructive algebra and systems theory, volume 53 of Verhandelingen, Afdeling Natuurkunde. Eerste Reeks. Koninklijke Nederlandse Akademie van Wetenschappen [Proceedings, Physics Section. Series 1. Royal Netherlands Academy of Arts and Sciences]. Royal Netherlands Academy of Arts and Sciences, Amsterdam, The Netherlands, 2006. ISBN 90-6984-477-X.

Hirota:1989:IRD

[HI89] Ryogo Hirota and Masaaki Ito. Introduction to REDUCE — Doing Symbolic Computation on PC. Science sha, Tokyo, Japan, June 1989. ISBN ???? ???? pp. LCCN ????

Hietarinta:1991:AFM

[Hie91] J. Hietarinta. From an analytical formula to a movie by way of REDUCE and C. In Proc. of the Workshop on Symbolic and Numeric Computation, pages 117–126. Research Reports, Computing Centre of Helsinki University, 1991.

Hearn:1983:RUM

[HM^{+}83] A. C. Hearn, J. D. Marti, et al. REDUCE user’s manual, version 3.0 (standard LISP report). Rand publication cp78 (4/83) (uucs-78-101)

(university of utah, technical report tr-6, 1978), Rand Corporation, Santa Monica, CA, USA, 1983. 1– 135 pp.

Harlander:2024:DCM

[HM24] Robert V. Harlander and Jean-Philippe Martinez. The development of computational methods for Feynman diagrams. European Physical Journal H, 49(1):4:1–4:45, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/
epjh/s13129-024-00067-6.

Hollenhorst:1992:DDF

[Hol92] Manfred Hollenhorst. The design of digital filters using REDUCE. SIG-SAM Bulletin, 26(1):10–12, January 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hehl:1996:RSC

[HPR96] F. W. (Friedrich W.) Hehl, R. A. (Roland A.) Puntigam, and H. (Hanns) Ruder, editors. Relativity and scientific computing: Computer algebra, numerics, visualization. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60361-1. LCCN QC173.55 .R445 1996.

Houstis:1992:AIE

[HR92] Elias N. Houstis and John R. Rice, editors. Artificial intelligence, expert systems, and symbolic computing: selected and revised papers from the IMACS 13th World Congress, Dublin, Ireland, July 1991. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Hearn:1983:RIG

[HS83] A. C. Hearn and L. R. Seward. REDUCE installation guide for the DEC VAX-11 series running UNIX, vers. 3.0. Report Rand Publication CP84 (4/83), The Rand Corporation, Santa Monica, CA, USA, April 1983.

Hearn:1995:CAS

[HS95] Anthony C. Hearn and Eberhard Schrüfer. A computer algebra system based on order-sorted algebra. Journal of Symbolic Computation, 19(1/2/3):65–79 (or 65– 77??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Hartley:1991:CIC

[HT91] David Hartley and Robin W. Tucker. A constructive implementation of the Cartan–Kähler theory of exterior differential systems. Journal of Symbolic Computation, 12(6): 655–668 (or 655–667??), December 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hartley:1995:GBC

[HT95] David Hartley and Philip Tuckey. Gröbner bases in Clifford and Grassmann algebras. Journal of Symbolic Computation, 20(2):197–206 (or 197–205??), August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hughes:1990:SCF

[Hug90] D. I. Hughes. Symbolic computation with Fermions. Journal of Symbolic Computation, 10(6):657–664, December 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Husberg:1981:PIR

[Hus81] N. Husberg. Preliminary II REDUCE-2 and Analitik-74, a comparison. Technical report, Helsinki University of Technology Computing Center, Helsinki, Finland, November 1981.

Hulshof:1983:SRF

[Hv83] B. J. A. Hulshof and J. A. van Hulzen. Some REDUCE facilities for pretty printing subscripts and formal derivatives. SIGSAM Bulletin, 17(1):16–20, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hulshof:1985:ECP

[Hv85] B. J. A. Hulshof and J. A. van Hulzen. An expression compression package for REDUCE based on factorization and controlled expansion. In Buchberger and Caviness [BC85], pages 315–316. CO-DEN LNCSD9. ISBN 0-387-15983-5 (v. 1), 0-387-15984-3 (v. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1985. Vol. 2 edited by: Bob F. Caviness. “Jointly organized by the ACM Special Interest Group on Symbolic and Algebraic Manipulation (SIGSAM) and by the Symbolic and Algebraic Manipulation Group in Europe (SAME)”–Vol. 2, pref. Contents: v. 1. Invited lectures — v. 2. Research contributions.

Hulshof:1983:R

[HvH^{+}83] B. Hulshof, A. van Hulzen, et al. REDUCE, 1983.

IEEE:1978:MPA

[IEE78] IEEE, editor. Micro 11: proceedings: the 11th annual Microprogramming Workshop, November 19–22, 1978, Asilomar Conference Ground, Pacific Grove, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978. ISSN 0194-1895. LCCN QA76.6 .W69a. Catalog number 78 CH1411-8.

Ito:1985:RPF

[IK85] Masaaki Ito and Fujio Kako. A REDUCE program for finding conserved densities of partial differential equations with uniform rank. Computer Physics Communications, 38(3):415–419, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ilyin:1991:SST

[IK91] V. A. Ilyin and A. P. Kryukov. Symbolic simplification of tensor expressions using symmetries, dummy indices and identities. In Watt [Wat91], pages 224–228. ISBN 0-89791-437-6. LCCN QA 76.95 I59

Ilyin:1996:APT

[IK96] V. A. Ilyin and A. P. Kryukov.

ATENSOR: REDUCE program for tensor simplification. Computer Physics Communications, 96(1):36– 52, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Igonin:2003:SCI

[IKK03] S. Igonin, P. H. M. Kersten, and I. Krasil′shchik. On symmetries and cohomological invariants of equations possessing flat representations. Differential Geom. Appl., 19(3):319–342, 2003. CO-DEN DGAPEO. ISSN 0926-

2245. URL http://diffiety.ac.
ru/preprint/2002/07_02abs.htm.

Ilyin:1989:FAC

[IKRT89a] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodioniov, and A. Yu. Taranov. Fast algorithm for calculation of Dirac’s gamma-matrices traces. SIGSAM Bulletin, 23(4):15–24, October 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ilyin:1989:HSD

[IKRT89b] V. A. Ilyin, A. P. Kryukov, A. Ya.

Rodioniov, and A. Yu. Taranov. High speed Dirac algebra calculations in a space of arbitrary dimension by means of a computer algebra system. Computer Physics Communications, 57(1-3):505–506, December 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ilyin:1991:PIF

[IKRT91] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodinov, and A. Yu. Taranov. PC implementation of fast Dirac matrix trace calculations. In Watt [Wat91], pages 456–457. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Inada:1980:FLS

[Ina80] Nobuyuki Inada. Fortran-based LISP system for REDUCE. Technical report, Information Science Laboratory, The Institute of Physical and Chemical Research, ????, 1980.

Ioakimidis:1990:CSI

[Ioa90a] N. I. Ioakimidis. Construction of singular integral equations for interacting straight cracks by using RE-DUCE. Technical report, Division of Applied Mathematics and Mechanics, School of Engineering, University of Patras, Patras, Greece, 1990.

Ioakimidis:1990:CEC

[Ioa90b] N. I. Ioakimidis. Construction of the equation of caustics in dynamic plane elasticity problems with the help of REDUCE. Technical report, Division of Applied Mathematics and Mechanics, School of Engineering, University of Patras, Patras, Greece, 1990.

Ioakimidis:1991:CSI

[Ioa91a] N. I. Ioakimidis. Construction of singular integral equations for interacting straight cracks by using Reduce. Engineering fracture mechanics, 40(6):1179–1184, ???? 1991. CODEN EFMEAH. ISSN 0013-7944.

Ioakimidis:1991:CEC

[Ioa91b] N. I. Ioakimidis. Construction of the equation of caustics in dynamic plane elasticity problems with the help of Reduce. Computers and Structures, 41(2):407–409, ???? 1991. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ito:1985:RPE

[Ito85] M. Ito. A REDUCE program for evaluating a Lax pair form. Computer Physics Communications, 34: 325–331, 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ito:1988:RPH

[Ito88] Masaaki Ito. A REDUCE program for Hirota’s bilinear operator and Wronskian operations. Computer Physics Communications, 50 (3):321–330, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ito:1994:SRP

[Ito94] M. Ito. SYMCD — a RE-DUCE package for finding symmetries and conserved densities of systems of nonlinear evolution equations. Computer Physics Communications, 79(3):547–??, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Jenks:1976:SPA

[Jen76] Richard D. Jenks, editor. Symsac ’76: proceedings of the 1976 ACM Symposium on Symbolic and Algebraic Computation, August 10– 12, 1976, Yorktown Heights, New York. Academic Press, New York, NY, USA, 1976. LCCN QA155.7.E4 .A15 1976. US$20.00.

Jerosolimski:1994:NMF

[JL94] M. Jerosolimski and L. Levacher. New method for fast calculation of Jacobian matrices: Automatic differentiation for power system simulation. IEEE Transactions on Power Systems, 9(2):700–706, May 1994. CODEN ITPSEG. ISSN 0885-8950 (print), 1558-0679 (electronic).

Kadlecsik:1993:TMP

[Kad93] J. Kadlecsik. Tensor manipulation package in REDUCE. In De Groot and Nadrchal [DN93], pages 362– 363. ISBN 981-02-1245-3. LCCN QC19.2.I53 1992.

Kadlecsik:1996:RCP

[Kad96] J. Kadlecsik. Ricci calculus package in REDUCE. Computer Physics Communications, 93(2-3):265–282, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Kalina:1982:RRL

[Kal82] Zdenêk Kalina. Running REDUCE and Lisp/360 interactively under a Time Sharing Option. SIGSAM Bulletin, 16(2):3, May 1982. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kampas:2005:TUR

[Kam05] F. J. Kampas. Tricks of using Reduce to solve Kuhn–Tucker equations. Mathematica Journal, 9(4):686–689, 2005. URL

www.mathematica-journal.com/
issue/v9i4/contents/Tricks9-
4/Tricks9-4_2.html. See correc-

tion [Par14].

Kanada:1975:IHA

[Kan75] Y. Kanada. Implementation of HLISP and algebraic manipulation language REDUCE 2. Report 75-01, University of Tokyo Information Science Lab, Tokyo, Japan, 1975.

Kazasov:1987:LTR

[Kaz87] C. Kazasov. Laplace transformations in REDUCE 3. In Proc. EU-ROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 132–133. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Kazasov:1989:LTR

[Kaz89] C. Kazasov. Laplace transformations in REDUCE 3. In Davenport [Dav89], pages 132–133. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

Kleene:1994:SBP

[KC94] S. J. Kleene and H. C. Cejtin. Solving buffering problems with Mathematica software. Analytical biochemistry, 222(2):310–314, November 1, 1994. CODEN ANBCA2. ISSN 0003-2697.

Keady:1991:GRM

[Kea91] G. Keady. GENTRANs from RE-DUCE and from MACSYMA. Technical Report II/13, University of Waikato Mathematics Department, Waikato, New Zealand, September 1991.

Keady:1992:FSP

[Kea92] G. Keady. Fortran subroutines produced from computer algebra systems: using GENTRANS from RE-DUCE and from MACSYMA. In Noye et al. [NBC92], pages 265–272. ISBN 0-86396-172-X. LCCN ????

Kempelmann:1981:RAF

[Kem81] Helmut Kempelmann. Recursive algorithm for the fast calculation of the limit of derivatives at points of indeterminateness. SIGSAM Bulletin, 15(4):10–11, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kennedy:1980:CUC

[Ken80] A. D. Kennedy. Comments on the use of computer symbolic algebra for theoretical physics. Surveys in High Energy Physics, 2(1–2):127– 155, November 1980. ISSN 0142-2413 (print), 1477-2892 (electronic).

Kessler:1979:PPM

[Kes79] R. Kessler. PMETA — pattern matching META/REDUCE. Report USCG Op. Note No. 40, Univ. of Utah, CS Dept., Salt Lake City, UT, USA, January 1979.

Kornyak:1987:RPC

[KF87] V. V. Kornyak and R. N. Fedorova. A REDUCE program to calculate determining equations of Lie-Baecklund symmetries of differential equations. Technical Report P11-87-19, J.I.N.R., Dubna, USSR, 1987.

Kaneko:1989:PFS

[KK89] Toshiaki Kaneko and Setsuya Kawabata. A preprocessor for Fortran source code produced by RE-DUCE. Computer Physics Communications, 55(2):141–147, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Krasilshchik:1992:DRO

[KK92] I. S. Krasil′shchik and P. H. M.

Kersten. Deformations and recursion operators for evolution equations. Memorandum 1104, University of Twente, Enschede, The Netherlands, 1992.

Kersten:1993:GFN

[KK93] P. H. M. Kersten and I. S.

Krasil′shchik. Graded Frölicher– Nijenhuis brackets and the theory of recursion operators for super differential equations. Memorandum 1147, Twente University, Enschede, The Netherlands, 1993. 56 pp.

Krasilshchik:1994:DRO

[KK94] I. S. Krasil′shchik and P. H. M. Kersten. Deformations and recursion operators for evolution equations. In Geometry in partial differential equations, pages 114–154. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994.

Kersten:1995:GFN

[KK95a] P. H. M. Kersten and I. S.

Krasil′shchik. Graded Frölicher– Nijenhuis brackets and the theory of recursion operators for super differential equations. In V. V. Lychagin, editor, The interplay between differential geometry and differential equations, volume 167 of Amer. Math. Soc. Transl. Ser. 2, pages 91–129. American Mathematical Society, Providence, RI, USA, 1995. ISBN 0-8218-0428-6.

Krasilshchik:1995:GDE

[KK95b] I. S. Krasil′shchik and P. H. M. Kersten. Graded differential equations and their deformations: a computational theory for recursion operators. Acta Applicandae Mathematicae, 41(1–3):167–191, 1995. CO-DEN AAMADV. ISSN 0167-8019 (print), 1572-9036 (electronic). Geometric and algebraic structures in differential equations.

Krasilshchik:2000:SRO

[KK00] I. S. (Iosif Semenovich) Krasil′shchik and P. H. M. Kersten. Symmetries and recursion operators for classical and supersymmetric differential equations, volume 507 of Mathematics and its applications. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6315-9 (hardcover). xvi + 384 pp. LCCN QA377 .K675 2000. URL http://www.

loc.gov/catdir/enhancements/
fy0822/00037545-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0822/00037545-t.html.

Kersten:2002:ROH

[KK02a] P. H. M. Kersten and I. S. Krasil′shchik. From recursion operators to Hamiltonian structures: the factorization method. Memorandum 1624, Faculty of Mathematical Sciences, University of Twente, Enschede, The Netherlands, 2002.

Kersten:2002:CIC

[KK02b] Paul Kersten and Joseph Krasil′shchik.

Complete integrability of the coupled KdV–mKdV system. In Lie groups, geometric structures and differential equations—one hundred years after Sophus Lie (Kyoto/Nara, 1999), volume 37 of Adv. Stud. Pure Math., pages 151–171. Math. Soc. Japan, Tokyo, Japan, 2002.

MR2391504

[KK06] P. H. M. Kersten and I. S.

Krasil′shchik. The Cartan covering and complete integrability of the KdV–mKdV system. In Hanzon and Hazewinkel [HH06], pages 251–265. ISBN 90-6984-477-X.

Kersten:2002:ESS

[KKV02] P. H. M. Kersten, I. S. Krasil′shchik, and A. M. Verbovetsky. An extensive study of the N = 1 supersymmetric KdV equation. Memorandum 1656, Faculty of Mathematical Sciences, University of Twente, Enschede, The Netherlands, 2002.

URL http://www.math.utwente.
nl/publications/2002/1656abs.
html.

Kersten:2003:SKE

[KKV03a] P. H. M. Kersten, I. S. Krasil′shchik, and A. M. Verbovetsky. The N = 1 supersymmetric KdV equation: (non)local Hamiltonian and symplectic structures, recursions, and hierarchies. arxiv.org, page 15,

                                        2003. URL http://arxiv.org/
                                        abs/nlin/0305026.

Kersten:2003:GAH

[KKV03b] P. H. M. Kersten, I. S. Krasil′shchik, and A. M. Verbovetsky. A geometric approach to Hamiltonian structures for evolution equations. In ????, editor, Proceedings of the Conference “Kolmogorov and Contemporary Mathematics”, Moscow, Russia, pages 815–816. ????, ????, 2003.

Kersten:2004:HOC

[KKV04a] P. Kersten, I. Krasil′shchik, and A. Verbovetsky. Hamiltonian operators and ℓ^{∗}-coverings. Journal of Geometry and Physics, 50(1–4): 273–302, 2004. CODEN JGPHE5. ISSN 0393-0440 (print), 1879-1662 (electronic). URL http://

                                        diffiety.ac.ru/preprint/2002/
                                        06_02abs.htm.

Kersten:2004:NLH

[KKV04b] P. Kersten, I. Krasil′shchik, and A. Verbovetsky. (Non)local Hamiltonian and symplectic structures, recursions and hierarchies: a new approach and applications to the N = 1 supersymmetric KdV equation. Journal of Physics A (Mathematical and General), 37(18):5003– 5019, 2004. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Kersten:2004:ICS

[KKV04c] P. Kersten, I. Krasil′shchik, and A. Verbovetsky. On the integrability conditions for some structures related to evolution differential equations. Acta Applicandae Mathematicae, 83(1-2):167–173, 2004. CODEN AAMADV. ISSN 0167-8019 (print), 1572-9036 (electronic).

Kersten:2004:MAE

[KKV04d] P. H. M. Kersten, I. S. Krasil′shchik, and A. M. Verbovetsky. The Monge–Ampère equation: Hamiltonian and symplectic structures, recursions, and hierarchies. Memorandum 1727, Faculty EEMCS, University of Twente, Enschede, The Netherlands, 2004. 75 pp.

URL http://www.math.utwente.
nl/publications/2004/1727.pdf.

Kersten:2004:NCG

[KKV04e] Paul Kersten, Iosif Krasil′shchik, and Alexander Verbovetsky. Nonlocal constructions in the geometry of PDE. In A. G. Nikitin, V. M. Boyko, R. O. Popovych, and I. A. Yehorchenko, editors, Symmetry in nonlinear mathematical physics. Part 1, 2, 3, volume 2 of Pr. Inst. Mat. Nats. Akad. Nauk Ukr. Mat. Zastos., 50, Part 1, pages 412–423. Natsīonal. Akad. Nauk Ukraïni, Īnst. Mat., Kiev, Ukraine, 2004. URL http:/

/www.imath.kiev.ua/~snmp2003/
Proceedings/krasilshchik.pdf.

Kersten:2006:GSD

[KKV06] Paul Kersten, Iosif Krasil′shchik, and Alexander Verbovetsky. A geometric study of the dispersionless Boussinesq type equation. Acta Applicandae Mathematicae, 90(1-2): 143–178, 2006. CODEN AAMADV. ISSN 0167-8019.

Kersten:2010:IKD

[KKVV10] P. H. M. Kersten, I. S. Krasilâshchik, A. M. Verbovetsky, and R. Vitolo.

Integrability of Kupershmidt deformations. Acta Applicandae Mathematicae, 109(1):75–86, 2010. CO-DEN AAMADV. ISSN 0167-8019 (print), 1572-9036 (electronic).

Kersten:2013:HSG

[KKVV13] P. H. M. Kersten, I. S. Krasil′shchik, A. M. Verbovetsky, and R. Vitolo. Hamiltonian structures for general PDEs. In Kruglikov et al. [KLS12], pages 187–198. ISBN 3-642-26933-8. LCCN ???? URL http://xxx.

                                        tau.ac.il/pdf/0812.4895.

Klimov:1993:SMG

[KLR93] D. M. Klimov, V. V. Leonov, and V. M. Rudenko. The study of motion for a gyroscope with gimbal suspension: obtaining the highest approximations for a drift of Magnus. Journal of Symbolic Computation, 15(1):73–78, January 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kruglikov:2013:DEG

[KLS12] Boris Kruglikov, Valentin Lychagin, and Eldar Straume, editors. Differential Equations — Geometry, Symmetries and Integrability: The Abel Symposium 2008. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-26933-8. LCCN ????

Kruse:1983:ARE

[KO83] Hans-Guenther Kruse and Karin Ohlsen. About the realization of an extended, but really interactive REDUCE by integration of a small editing and executing system. SIG-SAM Bulletin, 17(1):21–25, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kolbig:1985:EEC

[Köl85] K. S. Kölbig. Explicit evaluation of certain definite integrals involving powers of logarithms. Journal of Symbolic Computation, 1(1): 109–114, March 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Konieczny:1995:RIS

[Kon95] Janusz Konieczny. Reduced idempotents in the semigroup of Boolean matrices. Journal of Symbolic Computation, 20(4):471–482, October 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kotorynski:1986:SLF

[Kot86] W. P. Kotorynski. Steady laminar flow through a twisted pipe of elliptical cross-section. Computers and Fluids, 14(4):433–444, 1986. CO-DEN CPFLBI. ISSN 0045-7930.

Klerer:1968:ISE

[KR68] Melvin Klerer and Juris Reinfelds, editors. Interactive systems for experimental applied mathematics. Academic Press, New York, NY, USA, 1968. LCCN QA76.5.A1 A2 1967. ACM Symposium on Interactive Systems for Experimental Mathematics. Washington, DC, USA.

Kryukov:1985:DDS

[KR85a] A. P. Kryukov and A. Ya. Rodinov.

Dynamic-debugging system for the REDUCE programs. SIGSAM Bulletin, 19(2):34–37, May 1985. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kryukov:1985:IR

[KR85b] A. P. Kryukov and A. Ya. Rodinov. Interactive REDUCE. SIG-SAM Bulletin, 19(3):43–45, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kyrukov:1985:IR

[KR85c] A. P. Kyrukov and A. Ya. Rodinov. Interactive REDUCE. SIG-SAM Bulletin, 19(3):43–45, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kryukov:1986:URC

[KR86] A. P. Kryukov and A. Y. Rodinov. Usage of Reduce for computations of group-theoretical weight of Feynman diagrams in Non-Abelian gauge theories. In Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation, pages 91– 93. ????, ????, July 21–23, 1986.

Kryukov:1987:CAD

[KR87] A. P. Kryukov and A. Ya. Rodinov. CTS — algebraic debugging system for REDUCE programs. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 233–243. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Keady:1993:AIS

[KR93] G. Keady and M. G. Richardson. An application of IRENA to systems of nonlinear equations arising in equilibrium flows in networks. In Bronstein [Bro93], pages 311–320. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Kryukov:TBD:URC

[KRxx] A. P. Kryukov and A. Ya. Rodionov.

Usage of REDUCE for computations of group-theoretical weight of Feynman diagrams in Non-Abelian gauge theories. Technical report, Institute of Nuclear Physics, Moscow, USSR, 19xx.

Krasilshchik:2003:LES

[Kra03] I. S. Krasil′shchik. The long exact sequence of a covering: three applications. Preprint DIPS 6/2003, The Diffiety Institute, ????, 2003. URL http:/

/www.imath.kiev.ua/~snmp2003/
Proceedings/krasilshchik.pdf.

Kryukov:1986:CRA

[KRL86] A. P. Kryukov, A. Y. Rodinov, and G. L. Litvinov. Construction of rational approximations by means of REDUCE. In Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation, pages 31– 33. ????, ????, July 21–23, 1986.

Klimov:1993:SEN

[KRL93] D. M. Klimov, V. M. Rudenko, and V. V. Leonov. Symbolic evaluation in the nonlinear mechanical systems. In Bronstein [Bro93], pages 53–54. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Kryukov:1987:PCR

[KRR87] A. P. Kryukov, A. Ya. Rodinov, and V. A. Rostovtsev. Pattern compilation in REDUCE. Technical Report P11-87-302, J.I.N.R., Dubna, USSR, 1987.

Kryukov:1988:NPT

[KRR88] A. P. Kryukov, A. Ya. Rodinov, and V. A. Rostovtsev. New programming tools for computing substitution rules in REDUCE system. Technical Report P11-88-402, J.I.N.R., Dubna, USSR, 1988.

Kryukov:1984:ARL

[Kry84] A. P. Kryukov. An antitranslator of the RLISP language. SIGSAM Bulletin, 18(3):12–15, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Krynko:1986:DRE

[Kry86a] A. Krynko. Dialogue in REDUCE: Experience and development. In SYMSAC’86 — Proceedings of the 1986 Symposium on Symbolic and Algebraic Manipulation, pages 107– 109. ACM Press, Waterloo, Ontario, Canada, July 1986. ISBN 0-89791-199-7.

Kryukov:1986:DRE

[Kry86b] A. P. Kryukov. Dialogue in RE-DUCE: Experience and development. In Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation, pages 107–109. ????, ????, July 21–23, 1986.

Krasilshchik:2003:NHS

[KVK03] I. S. Krasil′shchik, A. M. Verbovetsky, and P. H. M. Kersten. Nonlocal Hamiltonian, symplectic and recursion structures for N = 1 supersymmetric KdV equation. In ????, editor, Proceedings of the Conference “Kolmogorov and Contemporary Mathematics”, Moscow, Russia, pages 817–818. ????, ????, 2003.

Lanyov:1994:RUC

[Lan94] A. V. Lanyov. REDUCE usage for calculations in chiral meson models. In Gruber and Tomassini [GT94], pages 179–181. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Liska:1987:EPF

[LD87] R. Liska and D. Drska. Evaluation of plasma fluid equations collision integrals using REDUCE. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, page 178. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Liska:1989:EPF

[LD89] R. Liska and L. Drska. Evaluation of plasma fluid equations collision integrals using REDUCE. In Davenport [Dav89], page 178. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

Liska:1990:FRP

[LD90] R. Liska and L. Drska. FIDE: A REDUCE package for automation of FInite difference method for solving pDE. In Watanabe and Nagata [WN90], pages 169–176. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Loos:1972:ATT

[Loo72] Rüdiger Loos. Analytic treatment of three similar Fredholm integral equations of the second kind with Reduce 2. SIGSAM Bulletin, ?? (??):32–40, January 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1977:ARN

[Loo77] Rüdiger Loos. Abstracts (RE-DUCE newsletter 1.1978). SIGSAM Bulletin, 11–12(4–1):86–88, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lloyd:1990:RBL

[LP90] N. G. Lloyd and J. M. Pearson. REDUCE and the bifurcation of limit cycles. Journal of Symbolic Computation, 9(2):215–224, February 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Leler:1985:IGI

[LS85] Wm Leler and Neil Soiffer. An interactive graphical interface for REDUCE. SIGSAM Bulletin, 19 (3):17–23, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Liska:1991:ADS

[LS91] R. Liska and M. Yu. Shashkov. Algorithms for difference schemes construction on non-orthogonal logically rectangular meshes. In Watt [Wat91], pages 419–426. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Lu:1996:SCD

[LS96] X. Y. Lu and S. K. Spurgeon. Symbolic computation for dynamic sliding mode controller design. In IEE, editor, IEE Colloquium on Symbolic Computation for Control (Digest No: 1996/078), 2 April 1996, pages 4/1–4/5. IEE, London, UK, 1996. ISBN ???? LCCN ????

MacCallum:1994:ACG

[M^{+}94] M. A. H. MacCallum et al., editors. Algebraic computing in general relativity: lecture notes from the First Brazilian School on Computer Algebra, Rio de Janeiro, Brazil, volume 2 of Lecture notes from the First Brazilian School on Computer Algebra. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853646-1. LCCN QC173.6.B73 1989. US$45.00.

MacCallum:1988:ODE

[Mac88] M. A. H. MacCallum. An ordinary differential equation solver for RE-DUCE. In Proc. of ISSAC ’88, volume 358, pages 196–205. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988.

MacCallum:1989:ODE

[Mac89] M. A. H. MacCallum. An ordinary differential equation solver for REDUCE. In Gianni [Gia89], pages 196–205. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Maguire:1981:PTR

[Mag81] Gerald Quentin Maguire, Jr. Program transformation in REDUCE using rule sequencing. Master’s thesis, Department of Computer Science, The University of Utah, Salt Lake City, UT, USA, March 1981.

Malm:1982:PRF

[Mal82] Bengt Malm. Program in Reduce for finding explicit solutions to certain ordinary differential equations. Lecture Notes in Computer Science, 144:289–293, 1982. CO-DEN LNCSD9. ISBN 3-540-11607-9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Man:1993:CCF

[Man93] Yiu-Kwong Man. Computing closed form solutions of first order ODEs using the Prelle–Singer procedure. Journal of Symbolic Computation, 16(5):423–444 (or 423– 443??), November 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Marti:1978:MRT

[Mar78] Jed Marti. The META/REDUCE translator writing system. ACM SIGPLAN Notices, 13:42–49, 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marti:1988:GIR

[Mar88] J. Marti. A graphics interface to REDUCE. In Proc. AAECC-6 1988, Lecture Notes in Computer Science, volume 357, pages 274– 296. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988.

Marti:1989:GIR

[Mar89] J. Marti. A graphics interface to REDUCE. In Mora [Mor89], pages 274–296. ISBN 3-540-51083-4. LCCN ????

Marti:1993:REA

[Mar93] Jed Marti. RLISP ’88: an evolutionary approach to program design and reuse, volume 42 of World Scientific series in computer science. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1479-0. xiii + 254 pp. LCCN QA76.73.L23 M36 1993.

Mathews:1989:USM

[Mat89] J. Mathews. Using a symbol manipulation program in statics to compute centroids and moments. CoED, 9(1):52–55, January–March 1989. CODEN CWLJDP. ISSN 0736-8607.

Mazzarella:1985:ISO

[Maz85] Giuseppe Mazzarella. Improved simplification of odd and even functions in REDUCE. SIGSAM Bulletin, 19(2):29–30, May 1985. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McCrea:1984:URF

[McC84] J. D. McCrea. The use of RE-DUCE in finding exact solutions of the quadratic Poincaré gauge field equations. In Classical General Relativity, pages 173–182. Cambridge University, 1984.

McCrea:1994:RGR

[McC94] J. D. McCrea. REDUCE in general relativity and Poincaré gauge theory. In MacCallum et al. [M^{+}94], pages 173–176. ISBN 0-19-853646-1. LCCN QC173.6.B73 1989. US$45.00.

Melenk:1974:FRT

[Mel74] H. Melenk. Formelmanipulation mit REDUCE auf dem TR 440. Technical report, Grossrechenzentrum für die Wissenschaft in Berlin, Berlin, Germany, May 1974.

Melenk:1993:ASS

[Mel93a] H. Melenk. Automatic symbolic solution of nonlinear equation systems in REDUCE. In G. Jacob, N. E. Oussous, and S. Steinberg, editors, Proceedings of the 1993 International IMACS Symposium on Symbolic Computation, pages 175–180. IMACS, Laboratoire d’Informatique Fondamentale de Lille, France, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1993.

Melenk:1993:ASN

[Mel93b] Herbert Melenk. Algebraic solution of nonlinear equation systems in REDUCE. Preprint TR 93-2, Konrad-Zuse-Zentrum für Informationstechnik, Berlin, Germany, January 1993.

Melenk:2014:RSM

[Mel14] Herbert Melenk. REDUCE symbolic mode primer. Report, Konrad-Zuse-Zentrum für Informationstechnik Berlin, Takustrasse 7, 14195 Berlin-Dahlem, Germany, May

2, 2014. URL http://reduce-
algebra.sourceforge.net/doc/
primer.html.

Marti:1983:RCM

[MF83] Jed Marti and John Fitch. RE-DUCE 2 for CP/M. SIGSAM Bulletin, 17(1):26–27, 37, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moritsugu:1988:PIF

[MG88] S. Moritsugu and E. Goto. A proposal for improvement of facilities of REDUCE. Technical report, Department of Information Science, University of Tokyo, Tokyo, Japan, December 1988.

Marti:1985:RLB

[MH85] Jed B. Marti and Anthony C.

Hearn. REDUCE as a Lisp benchmark. SIGSAM Bulletin, 19(3):8– 16, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moussiaux:1993:CRPa

[MM93] A. Moussiaux and R. Mairesse.

CONVODE: a REDUCE package for differential equations. In De Groot and Nadrchal [DN93], pages 134–143. ISBN 981-02-1245-3. LCCN QC19.2.I53 1992.

Melenk:1988:GBC

[MMN88] Herbert Melenk, H. M. Möller, and Winfried Neun. On gröbner bases computation on a supercomputer using REDUCE. Technical Report Preprint SC 88-2, Konrad-Zuse-Zentrum für Informationstechnik Berlin ZIB, Berlin, Germany, January 1988.

Melfo:1992:CCV

[MN92] A. Melfo and L. A. Núñez. Checking collineation vectors with REDUCE. General relativity and gravitation, 24(11):1125–1129, 1992. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). A package of programs for writing and checking the solutions to the equations for various types of collineations (symmetries of the metric, Christoffel, Riemann and Ricci tensors) is presented. Some examples of previously found collineations that have been checked are given, and new results reported.

Monagan:1992:HIT

[Mon92] Michael B. Monagan. A heuristic irreducibility test for univariate polynomials. Journal of Symbolic Computation, 13(1):47–58 (or 47–57??), January 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mora:1989:AAA

[Mor89] T. Mora, editor. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes. 6th International Conference, AAECC-6. Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51083-4. LCCN ????

Moussiaux:1993:CRPb

[Mou93] A. Moussiaux. CONVODE: a RE-DUCE package for solving differential equations. Journal of Computational and Applied Mathematics, 48 (1/2):157–166, October 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Seventh Spanish symposium on orthogonal polynomials and applications: September 1991, Granada, Spain.

Mohring:1979:SEA

[MS79] H. J. Mohring and A. Schiller. Some experience in applying the algebraic system Reduce to the calculation of scattering processes in QED and QCD. Technical report, ????, ????, 1979. ?? pp.

MacCallum:1991:ACR

[MW91a] Malcolm A. H. MacCallum and Francis J. Wright, editors. Algebraic computing with REDUCE: lecture notes from the First Brazilian School on Computer Algebra. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-853444-2, 0-19-853443-4 (paperback). LCCN QA155.7.E4 B73 1989. Prepared withLaTEX.

McCallum:1991:ACR

[MW91b] M. McCallum and F. Wright. Algebraic Computation with REDUCE. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN ???? ???? pp. LCCN ????

Noye:1992:CTA

[NBC92] John Noye, Basil Benjamin, and Len Colgan, editors. Computational techniques and applications: proceedings of 5th International Computational Techniques and Applications Conference, held at The University of Adelaide, 14–17 July, 1991, Computational Techniques and Applications. Australian Mathematics Society, Adelaide, South Australia, Australia, 1992. ISBN 0-86396-172-X. LCCN ????

Norman:1996:IRJ

[NF96] A. Norman and J. Fitch. Interfacing REDUCE to Java. In Calmet and Limongelli [CL96], pages 271–276. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996.

Ng:1979:SAC

[Ng79] Edward W. Ng, editor. Symbolic and algebraic computation: EU-ROSAM ’79, an International Symposium on Symbolic and Algebraic Manipulation, Marseille, France, June 1979, volume 72 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-09519-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E88 1979.

Ng:1989:CCU

[Ng89a] Tze Beng Ng. Computation of the cohomology of BŜO_{n}⟨16⟩ for 23 ≤ n ≤ 26 using REDUCE. Journal of Symbolic Computation, 7(1): 93–100 (or 93–99??), January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ng:1989:CCN

[Ng89b] Tze Beng Ng. Computation of the cohomology of BŜO_{n} < 16 > for 23 ≤ n ≤ 26 using REDUCE. Journal of Symbolic Computation, 7 (1):93–99, January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ng:1989:CCB

[Ng89c] Tze Beng Ng. Computation of the cohomology of BSO_{n} (16) for 23 ≤ n ≤ 26 using REDUCE. Journal of Symbolic Computation, 7(1):93–99, January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nakamura:1989:SCS

[NM89] Hideharu Nakamura and Shouichi Matsui. Symbolic Computation in Structural Mechanics using RE-DUCE. Gihodo Shuppan Company Ltd., 1-11-41, Akasaka, Minato-Ku, 107 Tokyo, Japan, 1989. ISBN ???? ???? pp. LCCN ????

Norman:1978:TRS

[Nor78] Arthur Norman. Towards a RE-DUCE solution to SIGSAM problem 7. SIGSAM Bulletin, 12(4): 14–18, November 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norton:1980:NAL

[Nor80] Lewis M. Norton. A note about Laplace transform tables for computer use. SIGSAM Bulletin, 14(2): 30–31, May 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norman:1990:CBI

[Nor90] A. C. Norman. A critical-pair/ completion based integration algorithm. In Watanabe and Nagata [WN90], pages 201–205. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Norman:1993:CDS

[Nor93] A. C. Norman. Compact delivery support for REDUCE. Lecture Notes in Computer Science, ??(722):331–340, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norman:1995:CDS

[Nor95] A. C. Norman. Compact delivery support for REDUCE. Journal of Symbolic Computation, 19(1/2/3): 133–143 (or 133–144??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Nieuwenhuis:1997:PBA

[NR97] R. Nieuwenhuis and A. Rubio. Paramodulation with built-in ACtheories and symbolic constraints. Journal of Symbolic Computation, 23(1):1–22 (or 1–21??), January 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Navarro-Saad:1985:AFT

[NSW85] Miguel Navarro-Saad and Kurt Bernardo Wolf. Applications of a factorization theorem for ninth-order aberration optics. Journal of Symbolic Computation, 1(2):235–240 (or 235–239??), June 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nucci:1990:IRP

[Nuc90] M. C. Nucci. Interactive RE-DUCE programs for calculating classical, non-classical and Lie– Bäcklund symmetries of differential equations. Preprint Math: 062090-051, Georgia Institute of Technology, School of Mathematics, Atlanta, GA, USA, 1990.

Nucci:1992:IRP

[Nuc92] M. C. Nucci. Interactive REDUCE programs for calculating classical, non-classical and approximate symmetries of differential equations. In William F. Ames and P. J. (Pieter Jacobus) Van der Houwen, editors, Computational and applied mathematics, II: differential equations: selected and revised papers from the IMACS World Congress, Dublin, Ireland, July 1991, pages 345–350. North-Holland, London, UK, 1992. ISBN 0-444-89702-X. LCCN QA370.C625 1991.

Nunez:1995:ACR

[Nun95] Luis A. Nunez. Algebraic computing with REDUCE. General relativity and gravitation, 27(4):457–??, 1995. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

Norman:2014:IR

[NV14] Arthur C. Norman and Raffaele Vitolo. Inside Reduce. Report, ????, ????, September 2014. vi + 65 pp. In-progress book manuscript kindly provided to the bibliographer on 14 September 2014.

Norman:1983:CVR

[NW83] Arthur C. Norman and Paul S. Wang. A comparison of the Vaxima and Reduce factorization packages. SIGSAM Bulletin, 17(1): 28–30, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ollongren:1988:PRF

[Oll88] Alexander Ollongren. On a particular restricted five-body problem: an analysis with computer algebra. Journal of Symbolic Computation, 6 (1):117–127 (or 117–126??), August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ochiai:1990:LAU

[ON90] Mitsuyuki Ochiai and Kiyokazu Nagatomo. Linear Algebra using RE-DUCE. Kindai Kagaku sha, Tokyo, Japan, January 1990. ISBN ???? ???? pp. LCCN ????

Ogilvie:1987:ASC

[OT87] J. F. Ogilvie and R. H. Tipping. On the analytic solution by computer algebra of some problems in the vibration-rotational spectroscopy of diatomic molecules. Journal of Symbolic Computation, 3(3):277– 281, June 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ozieblo:19xx:ARG

[Ozixx] A. Ozieblo. Application of RE-DUCE 2 in general theory of relativity. Technical report, Cyfronet, Krakow, Poland, 19xx.

Parreiras:2014:URC

[Par14] Sérgio O. Parreiras. Using Reduce to compute Nash equilibria: Classroom tools for game theory. Mathematica Journal, 16(??): ??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http://www.

mathematica-journal.com/2014/
03/using-reduce-to-compute-nash-

equilibria/. See [Kam05].

Padget:1990:UPS

[PB90] Julian Padget and Alan Barnes.

Univariate power series expansions in REDUCE. In S. Watanabe and Morio Nagata, editors, IS-SAC ’90 Proceedings of International Symposium on Symbolic and Algebraic Computation (Aug 20– 24 1990: Tokyo, Jpn), pages 82– 87. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-54892-5. LCCN ????

Petrov:2014:CPS

[PBG14] A. B. Petrov, R. Z. Bachtizin, and S. S. Ghots. Calculation of phonon spectra for one-dimensional chains using computer algebra system Reduce. In 2014 International Conference on Computer Technologies in Physical and Engineering Applications (ICCTPEA), pages 137– 138. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2014. ISBN 1-4799-5315-6.

Pearce:1981:AAO

[PH81] P. D. Pearce and R. J. Hicks.

The application of algebraic optimisation techniques to algebraic mode programs for REDUCE. SIG-SAM Bulletin, 15(4):15–22, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pearce:1983:DSE

[PH83] P. D. Pearce and R. J. Hicks. Data structures and execution times of Algebraic mode programs for RE-DUCE. SIGSAM Bulletin, 17(1): 31–37, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Podgorak:1984:ARC

[PR84] E. Podgórak and I. Romanowska. Application of REDUCE 2 to the construction of recurrence relations. Master’s thesis, Institute of Computer Science, University of Wroclaw, Wroclaw, Poland, 1984.

Passmore:2002:LPN

[PR02] Tim Passmore and A. J. Roberts. Low Prandtl number fluid convection modelled using symbolic algebra (REDUCE) and Matlab. The ANZIAM Journal, 44((C)):C590– C626, 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

Pasini:1991:SCO

[PSZ91] P. Pasini, F. Semeria, and C. Zannoni. Symbolic computation of orientational correlation function moments. Journal of Symbolic Computation, 12(2):221–231 (or 221– 232??), August 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rayna:1987:RSA

[Ray87] Gerhard Rayna. REDUCE— Software for Algebraic Computation. Symbolic Computation — Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96598-X (New York), 3-540-96598-X (Berlin). ix + 329 pp. LCCN QA155.7.E4 R39 1987.

Rayna:1990:JPK

[Ray90] G. Rayna. J. P. Keener on RE-DUCE software for algebraic computation. Bulletin of Mathematical Biology, 52(5):697–??, 1990. ISSN 0092-8240.

Roque:1991:CAS

[Rd91] Waldir L. Roque and Renato P. dos Santos. Computer algebra in spacetime embedding. Journal of Symbolic Computation, 12(3):381–389, September 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Renner:1991:NEE

[Ren91] F. Renner. Nonlinear evolution equations and the Painlevé analysis: a constructive approach with RE-DUCE. In Watt [Wat91], pages 289– 294. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Renner:1992:CRP

[Ren92] Friedrich Renner. A constructive REDUCE package based upon the Painlevé analysis of nonlinear evolutions equations in Hamiltonian and/or normal form. Computer Physics Communications, 70 (2):409–416, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Roelofs:1991:IMO

[RG91] M. Roelofs and P. K. H. Gragert. Implementation of multilinear operators in REDUCE and applications in mathematics. In Watt [Wat91], pages 390–396. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Richardson:1992:CTB

[Ric92] Daniel Richardson. Computing the topology of a bounded non-algebraic curve in the plane. Journal of Symbolic Computation, 14(6):619– 644 (or 619–643??), December 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rudenko:1991:ACA

[RLBS91] V. M. Rudenko, V. V. Leonov, A. F. Bragazin, and I. P. Shmyglevsky. Application of computer algebra to the investigation of the orbital satellite motion. In Watt [Wat91], pages 450–451. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Rodinov:1984:WNC

[Rod84] A. Ya. Rodinov. Work with non-commutative variables in the Reduce-2 system for analytical calculations. SIGSAM Bulletin, 18 (3):16–19, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rizzi:1985:USC

[RT85] Nicola Rizzi and Amabile Tatone. Using symbolic computation in buckling analysis. Journal of Symbolic Computation, 1(3):317–322 (or 317–321??), September 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ronveaux:1988:PSR

[RT88] A. Ronveaux and G. Thiry. Polynomial solution of recurrence relation and differential equation. SIGSAM Bulletin, 22(4):9–19, October 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ronveaux:1989:DES

[RT89] A. Ronveaux and G. Thiry. Differential equations of some orthogonal families in REDUCE. Journal of Symbolic Computation, 8(5): 537–541, November 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sage:1988:ATQ

[Sag88] Martin L. Sage. An algebraic treatment of quantum vibrations using REDUCE. Journal of Symbolic Computation, 5(3):377–384, June 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See erratum [Sag89].

Sage:1989:EAT

[Sag89] Martin L. Sage. Erratum: “An algebraic treatment of quantum vibrations using REDUCE” [J. Symbolic Comput. 5 (1988), no. 3, 377–384]. Journal of Symbolic Computation, 7 (1):101, 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Sag88].

Saito:1988:DUA

[SAK^{+}88] O. Saito, K. Abe, M. Kanno, S. Sobue, and T. Umeno. Development of user adaptive CAD system for control engineering using RE-DUCE. In Preprints of the 4th IFAC Symposium on Computer Aided Design in Control Systems — CADCS ’88, pages 333–338. ????, Beijing, P. R. China, August 23–25, 1988.

Sasaki:1979:APR

[Sas79] Tateaki Sasaki. An arbitrary precision real arithmetic package in RE-DUCE. In Ng [Ng79], pages 358– 368. ISBN 0-387-09519-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E88 1979.

Savage:1988:SCF

[Sav88] S. B. Savage. Symbolic computation of the flow of granular avalanches. American Society of Mechanical Engineers, Heat Transfer Division, (Publication) HTD, 105:57– 63, 1988. CODEN ASMHD8. ISSN 0272-5673.

Savage:1990:SCF

[Sav90] Stuart B. Savage. Symbolic computation of the flow of granular avalanches. Journal of Symbolic Computation, 9(4):515–530, April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schrufer:1982:IEC

[Sch82a] E. Schrüfer. An implementation of the exterior calculus in REDUCE: a status report. SIGSAM Bulletin, 16 (4):27–31, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1982:RPD

[Sch82b] F. Schwarz. A REDUCE package for determining Lie symmetries of ordinary and partial differential equations. Computer Physics Communications, 27:179–186, 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Schwarz:1983:RPS

[Sch83] Fritz Schwarz. A REDUCE package for series analysis by Hadamard’s theorem and QD schemes. SIG-SAM Bulletin, 17(1):38–44, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1985:ADS

[Sch85a] F. Schwarz. Automatically determining symmetries of partial differential equations. Computing (Vienna/New York), 34(2):91–106, 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1985:ADP

[Sch85b] Fritz Schwarz. An algorithm for determining polynomial first integrals of autonomous systems of ordinary differential equations. Journal of Symbolic Computation, 1(2): 229–234 (or 229–233??), June 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schwarz:1986:RPD

[Sch86] Fritz Schwarz. A REDUCE package for determining first integrals of autonomous systems of ordinary differential equations. Computer Physics Communications, 39 (2):285–296, February/March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Schrufer:1988:CNE

[Sch88a] Eberhard Schrüfer. A comment on “A note on Einstein metrics”. SIG-SAM Bulletin, 22(3):22–26, July 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1988:RPD

[Sch88b] F. Schwarz. A REDUCE package for determining first integrals of autonomous systems of ordinary differential equations. Computer Physics Communications, 39:285– 296, 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Schlegel:1991:DRS

[Sch91] H. Schlegel. Determination of the root system of semisimple Lie algebras from the Dynkin diagram. In Watt [Wat91], pages 239–240. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Seiler:1991:SRP

[Sei91] Werner M. Seiler. SUPERCALC — a REDUCE package for commutator calculations. Computer Physics Communications, 66(2-3):363–376, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Shablygin:1987:IEH

[Sha87] E. Shablygin. Integral equation with hidden eigenparameter solver: RE-DUCE and FORTRAN in tandem. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 186–191. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Schrufer:1987:ECC

[SHM87] Eberhard Schrüfer, Friedrich W.

Hehl, and J. Dermott McCrea. Exterior calculus on the computer: The REDUCE-package EXCALC applied to general relativity and to the Poincaré gauge theory. General relativity and gravitation, 19(2): 197–218, February 1987. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

Stauffer:1988:CSC

[SHWZ88] Dietrich Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky. Computer simulation and computer algebra: lectures for beginners. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-18909-2. ix + 155 pp. LCCN QA76.9.C65 C656 1988.

Stauffer:1989:CSC

[SHWZ89] Dietrich Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky. Computer simulation and computer algebra: lectures for beginners. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1989.

ISBN 0-387-51141-5. ix + 155 pp. LCCN QA76.9.C65 C656 1989.

Stauffer:1993:CSC

[SHWZ93] Dietrich Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky. Computer simulation and computer algebra: lectures for beginners. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1993. ISBN 3-540-56530-2 (Berlin), 0-387-56530-2 (New York). x + 387 pp. LCCN QA76.9.C65 C656 1993.

Suzuki:1987:FTS

[SI87] Hiromi Suzuki and Shinichi Iwamoto. Fast transient stability solution using algebraic processing language ‘Reduce’. Electrical Engineering in Japan (English translation of Denki Gakkai Ronbunshi), 107(4):51–59, July/August 1987. CODEN EEN-JAU. ISSN 0424-7760.

Simon:1992:SMS

[Sim92] Barry Simon. Symbolic math software: It’s not just for mainframes anymore. PC Magazine, 11(14): 405–??, August 1992. CODEN PCMGEP. ISSN 0888-8507.

Saito:1988:SMC

[SKA88] Osami Saito, Mikihito Kanno, and Kenichi Abe. Symbolic manipulation CAD of control engineering by using Reduce. International Journal of Control, 48(2):781–790, August 1988. CODEN IJCOAZ. ISSN 0020-7179.

Steeb:1992:ACR

[SL92] Willi-Hans Steeb and Dirk Lewien. Algorithms and computation with REDUCE. Bibliographisches Institut, Mannheim, Germany, 1992. ISBN 3-411-15651-1. viii + 158 pp. LCCN QA76.9.A43 S74 1992.

Socorro:1998:CAG

[SMH98] José Socorro, Alfredo Macías, and Friedrich W. Hehl. Computer algebra in gravity: Reduce–Excalc programs for (non-) Riemannian space-times. I. Computer Physics Communications, 115(2–3):264– 283, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S0010465598001337.

Saito:1989:DCA

[SMU^{+}89] O. Saito, Y. Miura, T. Umeno, T. Ueyama, and K. Abe. Development of computer aided design and analysis system for robot manipulators using algebraic manipulation language Reduce. IFAC Proceedings Series, 7:231–236, 1989. CODEN IPSEET. ISSN 0741-1146.

Snyder:1993:FAG

[Sny93] Wayne Snyder. A fast algorithm for generating reduced ground rewriting systems from a set of ground equations. Journal of Symbolic Computation, 15(4):415–450, April 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Soma:1985:RAR

[Som85] Takashi Soma. Recent applications of REDUCE in RIKEN. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 181–182. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1985.

Spiridonova:1987:SEA

[Spi87] M. Spiridonova. Some extensions and applications of REDUCE system. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 136–137. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Spiridonova:1989:SEA

[Spi89] M. Spiridonova. Some extensions and applications of REDUCE system. In Davenport [Dav89], pages 136–137. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

Smith:1983:IMM

[SS83] Paul Smith and Leon Sterling. Of integration by man and machine. SIGSAM Bulletin, 17(3–4):21–24, August/November 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Surguladze:1991:APC

[SS91] L. R. Surguladze and M. A. Samuel. Algebraic perturbative calculations in high energy physics. Methods, algorithms, computer programs and physical applications. In Watt [Wat91], pages 439–447. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Suppes:1989:ICT

[ST89a] Patrick Suppes and Shuzo Takahashi. An interactive calculus theorem-prover for continuity properties. Journal of Symbolic Computation, 7(6):573–590, June 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Surguladze:1989:LPM

[ST89b] L. R. Surguladze and F. V. Tkachov.

LOOPS: Procedures for multiloop calculations in quantum field theory for the REDUCE system. Computer Physics Communications, 55 (2):205–215, September 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Steeb:1994:QMU

[Ste94] Willi-Hans Steeb. Quantum mechanics using computer algebra: includes sample programs for RE-DUCE, MAPLE, MATHEMATICA and C++. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-1770-6. viii + 189 pp. LCCN QC174.17.D37 S74 1994.

Stoutemyer:1977:AEA

[Sto77] David R. Stoutemyer. Automatic error analysis using computer algebraic manipulation. ACM Transactions on Mathematical Software, 3(1):26–43, March 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sarlet:1992:RPS

[SV92] W. Sarlet and J. Vanden Bonne. REDUCE-procedures for the study of adjoint symmetries of secondorder differential equations. Journal of Symbolic Computation, 13 (6):683–693, June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tao:1990:SAM

[Tao90] Qingsheng Tao. Symbolic and algebraic manipulation for formulae of interpolation and quadrature. In Watanabe and Nagata [WN90], page 306. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Thas:1989:CRMa

[Tha89a] C. Thas. A collection of REDUCE and MACSYMA programs about college geometry. part 1. Technical Report 5, State University of Gent, Gent, Belgium, September 1989.

Thas:1989:CRMb

[Tha89b] C. Thas. A collection of REDUCE and MACSYMA programs about college geometry. part 2. Technical Report 5, State University of Gent, Gent, Belgium, September 1989.

Triulzi:2000:OSU

[Tri00] Arrigo Triulzi. OpenMath support under CSL-hosted Reduce. SIG-SAM Bulletin, 34(2):27–30, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Trenkov:1991:ARS

[TSD91] I. Trenkov, M. Spiridonova, and M. Daskalova. An application of the REDUCE system for solving a mathematical geodesy problem. In Watt [Wat91], pages 448–449. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Umeno:1987:SPC

[UAYS87] Takaji Umeno, Kenichi Abe, Syuichi Yamashita, and Osami Saito. Software package for control design based on the algebraic theory using symbolic manipulation language Reduce. In ????, pages 1102–1106. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. IEEE Service Cent. Piscataway, NJ, USA.

Ueberberg:1992:ECR

[Ueb92] Johannes Ueberberg. Einführung in die Computeralgebra mit RE-DUCE. (German) [Introduction to computer algebra with REDUCE]. Bibliographisches Institut, Mannheim, Germany, 1992. ISBN 3-

411-15781-X.  336 pp.  URL

http://www.zentralblatt-math.
org/zmath/en/search/?an=0783.

68005. Für Mathematiker, Informatiker und Physiker. [For mathematicians, information scientists and physicists].

Ucoluk:1982:PER

[ÜHK82] G. Üçoluk, A. Hacınlıyan, and E. Karabudak. A proposal for extensions to REDUCE. SIGSAM Bulletin, 16(2):4–14, May 1982. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Urintsev:1991:CRA

[US91] A. L. Urintsev and A. V. Samoilov. Complex reduce-programs for analytic solution of some problems of beam transport systems. In D. V. Shirkov, V. A. Rostovtsev, and V. P. Gerdt, editors, In: 4th International Conference on Computer Algebra in Physical Research, pages 438–442. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1991.

Umeno:1989:SCA

[UYSA89] Takaji Umeno, Syuichi Yamashita, Osami Saito, and Kenichi Abe. Symbolic computation application for the design of linear multivariable control systems. Journal of Symbolic Computation, 8(6):581– 588, December 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vandenHeuvel:1986:ASR

[van86a] Pim van den Heuvel. Adding statements to REDUCE. SIGSAM Bulletin, 20(1 and 2):8–14, February/ May 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vandenHeuvel:1986:SER

[van86b] Pim van den Heuvel. Some experiments in REDUCE related to the calculation of Gröbner bases. Technical report, Twente University of Technology, Department of Computer Science, Enschede, The Netherlands, June 1986.

vanHulzen:1988:FMM

[van88] J. A. van Hulzen. Formule manipulatie m.b.v. REDUCE (in Dutch). Technical report, Twente University of Technology, Department of Computer Science, Enschede, The Netherlands, October 1988.

vanHulzen:19xx:SSC

[vanxx] J. A. van Hulzen. SCOPE 1, a source-code optimization package for REDUCE — user’s manual. Technical report, Twente University of Technology, Department of Computer Science, Enschede, The Netherlands, 19xx.

Vega:1991:WRA

[Veg91] Laureano González Vega. Working with real algebraic plane curves in REDUCE: the GCUR package. In Stephen M. Watt, editor, Proc. of the 1991 International Symposium on Symbolic and Algebraic Computation, pages 397–402. ACM Press, New York, NY 10036, USA, July 1991.

vanHulzen:1982:EAP

[vH82] J. A. van Hulzen and B. J. A. Hulshof. An expression analysis package for REDUCE. SIGSAM Bulletin, 16(4):32–44, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanHulzen:1989:COP

[vHGv89] J. A. van Hulzen, B. J. A. Hulshof, B. L. Gates, and M. C. van Heerwaarden. A code optimization package for REDUCE. In Gonnet [Gon89], pages 163–170. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

vandenHeuvel:1987:SSP

[vHv87] P. van den Heuvel, B. J. A. Hulshof, and J. A. van Hulzen. Some simple Pretty-Print facilities for RE-DUCE. SIGSAM Bulletin, 21(1): 14–17, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Visconti:1973:TIC

[Vis73] A. Visconti, editor. Third International Colloquium on Advanced Computing Methods in Theoretical Physics: Marseille, June 25–29, 1973. Centre de physique theorique, C.N.R.S., Marseille, France, 1973. ISBN ???? LCCN ???? Two volumes.

Vinitsky:1987:URS

[VR87] S. I. Vinitsky and V. A. Rostovtsev. A use of REDUCE system in problems of hydrogen atom in an electric field. Preprint P11-87-303, J.I.N.R., Dubna, USSR, 1987.

Vulcanov:2002:GTD

[Vul02] Dumitru N. Vulcanov. Gravity, torsion, Dirac field and computer algebra using MAPLE and REDUCE. arxiv.org, page 20, September 25,

                                        2002. URL http://arxiv.org/
                                        abs/gr-qc/0209096.

Vulcanov:2003:CDE

[Vul03] Dumitru N. Vulcanov. Calculation of the Dirac equation in curved spacetimes with possible torsion using MAPLE and RE-DUCE. Computer Physics Communications, 154(3):205–218, 2003.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

vanHeerwaarden:1988:PPF

[vv88] M. C. van Heerwaarden and J. A.

van Hulzen. Pretty print facilities for REDUCE. Memorandum INF-88-36, Twente University of Technology, Department of Computer Science, Enschede, The Netherlands, August 1988.

Verbovetsky:2012:ISG

[VVKK12] A. M. Verbovetsky, R. Vitolo, P. Kersten, and I. S. Krasil’shchik. Integrable structures for a generalized Monge–Ampère equation. Teoreticheskaya i Matematicheskaya Fizika, 171(2):208–224, 2012. CO-DEN TMFZAL. ISSN 0564-6162.

Wanas:1985:MPI

[Wan85] M. I. Wanas. Manipulation of parameters indicating the physical significance of any absolute parallelism space using REDUCE 2. In Tenth International Congress for Statistics, Computer Science, Social and Demographic Research. ????, ????, 1985.

Watt:1991:IPI

[Wat91] Stephen M. Watt, editor. ISSAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15–17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Wright:1985:EIA

[WD85] F. J. Wright and G. Dangelmayr.

Explicit iterative algorithms to reduce a univariate catastrophe to normal form. Computing: Archiv für informatik und numerik, 35(1): 73–83, 1985. CODEN CMPTA2. ISSN 0010-485X.

Winkelmann:1989:RBS

[WH89] Volker Winkelmann and Friedrich W. Hehl. REDUCE for beginners. six lectures on the application of computer algebra. In D. Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky, editors, Computer Simulation and Computer Algebra. Lectures for Beginners, chapter 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2nd edition, 1989.

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Weber:1979:PSR

[WR79] Lawrence A. Weber and Gerhard Rayna. Problem #11 solved in RE-DUCE: a case study in program translation. SIGSAM Bulletin, 13 (4):21–24, November 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Yamamoto:1987:RIP

[YA87] T. Yamamoto and Y. Aoki. RE-DUCE 3.2 on iAPX 86/286-based personal computers. In Proc. EU-ROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 134–135. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Yamamoto:1989:RIP

[YA89] T. Yamamoto and Y. Aoki. RE-DUCE 3.2 iAPX86/286-based personal computers. In Davenport [Dav89], pages 134–135. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

Yamamoto:1990:ACC

[YK90] Ichiro Yamamoto and Akira Kanagawa. Approximate column constants expressed in terms of design parameters for hot-wire thermal diffusion column. Journal of Nuclear Science and Technology, 27(1):49– 55, January 1990. CODEN JN-STAX. ISSN 0022-3131.

Yamartino:1991:ACA

[YP91] Robert J. Yamartino and Richard Pavelle. An application of computer algebra to a problem in stratified fluid flow. Journal of Symbolic Computation, 12(6):669–672, December 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zharkov:1993:CCI

[Zha93] A. Yu. Zharkov. Computer classification of the integrable coupled Kdv-like systems with unit main matrix. Journal of Symbolic Computation, 15(1):85–90, January 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:19xx:RNL

[ZZZxx] REDUCE network library, bibli-

                                        ography, 19xx. reduce-library@

rand.org, wird laufend ergänzt.